### v2

# VoCoT: Unleashing Visually Grounded Multi-Step Reasoning in Large Multi-Modal Models

Abstract
While large multi-modal models (LMMs) have exhibited impressive capabilities across diverse tasks, their effectiveness in handling complex tasks has been limited by the prevailing single-step reasoning paradigm. To this end, this paper proposes VoCoT, a multi-step Visually grounded object-centric Chain-of-Thought reasoning framework tailored for inference with LMMs. VoCoT is characterized by two key features: (1) object-centric reasoning paths that revolve around cross-modal shared object-level information, and (2) visually grounded representation of object concepts in a multi-modal interleaved and aligned manner, which effectively bridges the modality gap within LMMs during long-term generation. Additionally, we construct an instruction dataset to facilitate LMMs in adapting to reasoning with VoCoT. By introducing VoCoT into the prevalent open-source LMM architecture, we introduce VolCano. With only 7B parameters and limited input resolution, VolCano demonstrates excellent performance across various scenarios, surpassing SOTA models, including GPT-4V, in tasks requiring complex reasoning. Our code, data and model will be available at https://github.com/RupertLuo/VoCoT.
11footnotetext:  Co-first authorship.  Corresponding author.
1 Introduction
In recent years, the success of large language models (LLMs) [36; 37; 45] has been gradually extended to the multi-modal domain. By equipping LLM backbones [45; 46; 8; 16] with visual encoders [40] and efficient cross-modal alignment through generative training on image-text data [31; 42; 4; 63], constructed large multi-modal models (LMMs) possess the capabilities to perceive visual signals and engage in dialogue with users in multi-modal contexts [22; 62; 31; 58; 9; 2].
Despite the potential demonstrated by LMMs to serve as unified and versatile foundations for multi-modal information processing, even models like GPT-4V struggle in compositional tasks requiring complex reasoning [55; 53], such as spatial reasoning [10]. This phenomenon indicates a major limitation of current LMMs: the prevailing single-step question-to-answer (Q2A) inference paradigm that directly generates answers based on questions [9; 29]. As illustrated in Figure 1, correctly answering the question relies on analyzing the actions and relationships of multiple objects and reasoning step-by-step, which is almost impossible to accomplish in a single-step prediction. Moreover, the Q2A paradigm obscures the problem-solving process, limiting the interpretability of the LMM outputs. Conversely, the chain-of-thought (CoT) paradigm, which involves multi-step reasoning, has been widely explored in LLMs [19; 51; 49]. The effectiveness of textual CoTs suggests a promising direction for enhancing LMMs.
Figure 1: An example to compare different inference paradigms in LMMs. (a) A visual question that requires complex reasoning. (b) The conceptual object-centric reasoning path constructed to solve the problem. (c) Outputs of GPT-4V and the proposed VolCano. Hallucination is included in the output of GPT-4V. VoCalno performs multi-step reasoning in VoCoT format. Key objects are highlighted and colors indicate the correspondence between object descriptions and the grounded objects in the image. “box” represents the bounding box coordinates of mentioned objects.
However, in complex contexts where multi-modal information coexists, constructing effective multi-step reasoning paths faces several challenges: (1) Difficulty in integrating reasoning anchors within multi-modal contexts. Textual CoTs mainly extract key information from contexts, such as entities, as anchors and conduct multi-step reasoning around these anchors [57]. In multi-modal contexts, the anchor information is further required to be concepts shared between images and texts and establish connections between modalities. Existing work attempts to introduce anchor information in images, such as segmentation maps [54] and dot matrices [20], asking LMMs to utilize this information during inference. However, these methods require strong capabilities of LMMs to understand these information, thus are almost exclusively applicable to GPT-4V. (2) Limited grounding capabilities of LMMs. During the generation process, LMMs may fail to ground textual descriptions to the corresponding visual information, resulting in erroneous information generated. For example, GPT-4V incorrectly ground the target person to the waiter in Figure 1. The risk of producing hallucination [24; 48] hinders the model from performing effective multi-step reasoning.
To address these challenges, we introduce a framework to empower LMMs for effective and reliable multi-step reasoning. We propose VoCoT: Visually grounded object-centric Chain of Thought, a CoT format that is compatible with LMM inference: (1) As illustrated in Figure 1 (a, b), objects serve as fundamental semantic units in both images and text, effectively bridging multi-modal information. Therefore, VoCoT leverages objects as anchors for reasoning. LMMs are encouraged to conduct multi-step analysis on the properties of key objects, as well as the relationships between them, ultimately reaching a conclusion. (2) To ensure the reliability of information in reasoning paths, VoCoT represents each object in a visually grounded format: a tuple of <textual description, coordinates, corresponding visual representations>. Models are required to explicitly ground objects in the image by generating coordinates for them. Visual representations of objects are supplemented to enhance the cross-modal relevance in reasoning paths. This design mimics the reasoning habit of human, where we continuously reference the visual information of an object in the image when we mention it. (3) We propose a RefBind mechanism to efficiently obtain the visual representations of objects without additional computation. Specifically, RefBind indexes the representation of each object from the complete image representation based on its coordinate. Generally, VoCoT constructs multi-modal interleaved reasoning paths where cross-modal aligned anchors are incorporated.
Nevertheless, there is a significant disparity between VoCoT and the formats of existing visual instruction data. To address this, we further construct a dataset, VoCoT-Instruct-80K, to train LMMs for reasoning in the format of VoCoT. VoCoT-Instruct-80K is built on multiple data sources: (1) Reformulating structured reasoning paths from GQA [15]. (2) Supplementing visual question-answer pairs with thought processes. (3) Constructing complex questions and reasoning paths from images annotated with objects. By curating a wide range of data and leveraging assistance of GPT-4V, the presented dataset maintains both diversity and consistency in the desired format.
Based on the introduced VoCoT framework and dataset, we develop VolCano, a Visually grounded multi-modal Chain-of-thought reasoning model. With only 7B parameters and $336\times 336$ input resolution, VolCano surpasses GPT-4V on various benchmarks that require complex reasoning and compositional capabilities.
2 VoCoT: Visually grounded Object-centric Chain-of-Thoughts
In this section, we explain how to enable LMM to perform multi-step reasoning in the format of VoCoT. In Section 2.1, we elaborate on the formulation of VoCoT. Section 2.2 presents how to utilize existing data resources to construct instruction-tuning datasets that conform to the VoCoT format.
2.1 Formulating VoCoT
VoCoT requires LMMs to perform step-by-step reasoning based on the provided context. Following textual CoTs [19; 51; 49], the reasoning logic in VoCoT is primarily expressed in text but is not limited to specific formats. However, there exists a significant gap between multi-modal and text-only contexts. In order to construct effective and reliable reasoning paths in multi-modal contexts, we characterize VoCoT with two features: (1) Object-centric. Objects are the basic semantic units in images and can serve as anchors to establish connections between multi-modal contextual information. Therefore, VoCoTs are required to include important objects, followed by relevant information extraction and analysis. (2) Visually grounded. Key objects included in VoCoT should be represented by tuples of“<text description, coordinates, visual object representation>”, coordinates ground objects in the images while the visual representation enhances the cross-modal relevance in the reasoning paths.
2.2 VoCoT Dataset Construction
Table 1: Examples to illustrate the construction of VoCoT-formatted data from three data sources. Type 1 data are obtained by rules, while Type 2 and Type 3 data are obtained by leveraging GPT-4V.
Type 1 (GQA Source) Rule Input
[Object Info]: [Uncaptioned image]
Shelf: [0.224, 0.219, 0.386, 0.592], Door: [0.394, 0.176, 0.524, 0.645]
[Question]: What is common to the shelf and the door?
[SemanticStr]: select: shelf -> select: door -> common: [0, 1]
[FullAnswer]: The material, both the shelf and the door are wooden.
[Answer]: material.
Type 1 (GQA Source) Response
[Thought]: Find the shelf [0.224, 0.219, 0.386, 0.592]. Find the door [0.394, 0.176, 0.524, 0.645] . The question ask the common attribute of the two objects. The material, both shelf [0.224, 0.219, 0.386, 0.592] and door [0.394, 0.176, 0.524, 0.645] are wooden. So answer is material.
Type 2 (VQA-Based Source) GPT-4V Input
[Object Info]: TV: [0.78, 0.84, 0.97, 0.98], , Tie: [0.72, 0.12, 0.90, 0.60], [Uncaptioned image]
Toothbrush: [0.15, 0.37, 0.56, 0.49], Person: [0.0, 0.13, 0.81, 1.0]
[Question]: What might be the purpose behind the woman’s action?
[Answer]: The purpose behind her actions may be for entertainment.
Type 2 (VQA-Based Source) Response
[Thought]: The woman in the image [0.0, 0.13, 0.81, 1.0] is engaging in an unusual activity by using a toothbrush [0.15, 0.37, 0.56, 0.49] that is significantly larger than a standard one and wearing a tie [0.38, 0.6, 0.59, 0.99] that also appears to be oversized. … It’s plausible that the woman’s actions are meant to entertain or educate an audience.
Type 3 (Image-Only Source) GPT-4V Input
[Object Info]: Ferris wheel: [0.09, 0.40, 0.14, 0.54] , [Uncaptioned image]
Building: [0.40, 0.36, 0.58, 0.53], Statue: [0.72, 0.12, 0.90, 0.60],
Dock: [0.00, 0.56, 0.33, 0.67], Stair: [0.01, 0.69, 1.00, 1.00],
Type 3 (Image-Only Source) Response
[Question]: Where is this location?
[Refering of Thought]: First, The statue [0.72, 0.12, 0.90, 0.60] is
Merlion Statue, a mythical creature with the head of a lion and
the body of a fish, which is a mascot of Singapore. Secondly, the building [0.40, 0.36, 0.58, 0.53] … Based on the distinctive architecture of the landmarks, this image is taken in Singapore.
[Answer]: Singapore.
The Community has witnessed a surge in multi-modal instruction-following datasets, such as LLaVA-Instruct-150K[31], Valley-Instruct-65K[35] and ALLaVA [4]. However, none of these datasets meet the requirements of the VoCoT format, which includes: (1) responses to human instructions with CoT-formatted multi-step reasoning processes, (2) responses with visually grounded object-centric information, i.e., objects with corresponding coordinates. In this section, we introduce the pipeline to construct a VoCoT-formatted dataset from three types of data sources as summarized below.
Type 1: GQA Source
GQA [15] is a VQA dataset that includes structured information: each image is paired with a scene graph, and a SQL-like reasoning path over the corresponding scene graph is provided for each VQA pair. An example is exhibited in the first part of Table 1. Inspired by Shikra [6], we convert the SQL-like query statements (“[SementicStr]") and answers (“[FullAnswer]") into fluent and complete textual reasoning thoughts with a rule-based approach. Object bounding boxes from the GQA source are also utilized to supplement the anchor object information that appears in the verbalized reasoning thoughts.
Type 2: VQA-Based Source
Another intuitive way to construct data in VoCoT format is to supplement VQA data with multi-step reasoning processes in the middle of the question-to-answer process. With the assistance of GPT-4V, reasoning thoughts are generated based on images, questions, answers, and object information within the images. The second part of Table 1 provides an example. Furthermore, we control the output format through in-context learning. Specifically, a crafted sample is included in the input context. As overly simple questions may not require complex reasoning, we sample a subset of data from complex reasoning problems in LLaVA-Instruct [30] as the source.
Type 3: Image-Only Source
Although the aforementioned two construction methods are effective, the generated data is limited to existing questions. To enhance the richness of questions and reasoning logic, we leverage the powerful generative capabilities of GPT-4V to expand the constructed dataset. As shown in the third part of Table 1, GPT-4V is provided with images and object information and is prompted to generate complex questions, along with VoCoT-formatted reasoning paths and answers. In-context samples are also incorporated to ensure the correct output format. We choose LVIS [12] as the data source due to the diversity of objects included.
Ultimately, we present VoCoT-Instruct-80K, comprising 72K, 6K, and 2K samples constructed from Type 1, Type 2, and Type 3 sources, respectively. More details about the dataset construction process, including the rule-based conversion approach, prompts for GPT-4V, and quality control methods used are provided in Appendix A.1.
3 VolCano: A VoCoT-enhanced LMM
In this section, we introduce how to adapt a modern LMM to utilize the VoCoT framework. We present the architecture of VolCano in Section 3.1, and detail the model training process in Section 3.2.
3.1 Architecture
Figure 2: Illustration of the VolCano framework. The input and output are shown below and above the model, respectively. The blue and green rounded rectangles represent textual and visual tokens, respectively. Special tokens “[c]” and “[/c]” denotes the beginning and end of the coordinates (“[coor.]” in the figure). Coordinates are represented in text. In the output, we visualize coordinates by drawing corresponding boxes in the image for a better illustration. RefBind computes the visual representations of objects with the image features and predicted coordinates.
The overall model architecture of VolCano is shown in Figure 2. VolCano is built on top of a causal decoder-only LLM as the backbone. We incorporate a vision transformer (ViT) as the visual encoder to encode image inputs. A two-layer MLP is adopted as the connection module to map the output of the visual encoder into the input space of the language backbone.
Representations of Multi-modal Sequences
VolCano represents image-text data as an interleaved sequence of visual and textual tokens. Text inputs are tokenized and represented using the token embedding layer. Images and objects can appear at any position in the sequence and are represented by visual tokens. Images are encoded by the visual encoder. The resulting 2D feature maps are flattened into a 1D sequence of visual tokens. The connection module is applied to map the visual tokens to the same dimension as the embedded textual tokens.
Following the configuration of VoCoT in Section 2.1, each object is represented by a visually grounded format: “{textual description} [c] {coordinates} [/c] {visual representation}”, e.g., “dog [c] 0.27, 0.08, 0.92, 0.81 [/c] $V_{\textrm{dog}}$”. “[c]” and “[/c]” are special tokens denoting the beginning and ending of coordinates. We use bounding boxes $[x_{\textrm{min}},y_{\textrm{min}},x_{\textrm{max}},y_{\textrm{max}}]$ as coordinates of objects, $x$ and $y$ are normalized between 0 and 1 with respect to the size of the image. Coordinates are treated as text, undergoing tokenization and embedding.
In addition to text and coordinates, visual tokens of objects, such as $V_{\textrm{dog}}$, are supplemented to help the model reference the corresponding visual information in images. The visual tokens are obtained based on the coordinates and image tokens through the RefBind mechanism. The details of RefBind are explained in the next paragraph. Once the end of coordinates token, “[/c]”, is detected in the input or generated sequence, RefBind is activated to obtain the visual object tokens based on the coordinate between “[c]” and “[/c]”. The obtained object tokens are further appended after the “[/c]” token.
Figure 3: Illustration of the RefBind mechanism.
RefBind Mechanism
A straightforward method for representing objects is to crop the corresponding region in the image and encode each sub-image with the visual encoder. However, this method introduces additional computational costs and loses the contextual information of the complete image. For regions with very few pixels, representing objects with sub-images would introduce redundant information. To tackle with above issues, we propose the RefBind mechanism.
RefBind (short for “Reffering Bind”) is conceptually illustrated in Figure 3. Inspired by the RoI-align method in Faster-RCNN [41], given a bounding box and the encoded 2D grid features of the entire image, we can efficiently index the patches in which the target object appears. The features of these patches are flattened into a sequence that represents the object. RefBind relies solely on indexing operations without additional computation. Additionally, the object representation sequence obtained by RefBind inherently preserves contextual information of the whole image.
3.2 Training
The training of VolCano undergoes three stages summarized below. For more details on the utilized and the implementation settings, please refer to Section 4.1 and Appendix A.3.
Stage 1: Alignment Pre-training
The first stage aims to align visual representations with the LLM backbone. We utilize image-caption pairs from LLaVA-Pretrain [31; 29]. Both the visual encoder and the LLM backbone are frozen while the parameters in the connection module are updated.
Stage 2: Multi-modal Interleaved Pre-training
Following the alignment stage, the model is trained to adapt to multi-modal interleaved sequences and visually grounded object representations with RefBind. Three types of sequences are considered: (1) Image-caption pair constitutes the simplest form of multi-modal sequences. Different from the first stage, we utilize captions collected by ALLaVA [4] which provide detailed descriptions. (2) Multi-modal document includes multiple image-text pairs in a sequence. Based on the relevance between images and text, we filter a subset of documents from MMC4 [63]. (3) Grounded image caption further annotates the coordinates of objects in the caption. The representations of included objects are extended to the visually grounded format consistent with VoCoT. Flickr30K Entities [39] and a filtered subset of GRIT [38] are adopted. Besides the connection module, the language backbone is trained to model the multi-modal sequences.
Stage 3: Instruction Tuning
Based on the pre-trained model, the model is fine-tuned to follow instructions in multi-modal context and perform multi-step reasoning with VoCoT. We supplement the constructed VoCoT-Instruct-80K introduced in Section 2.2 and referring expression data [18; 6] to existing visual instruction data [29], which does not include CoT-form data. Similar to the last stage, both the language backbone and the connection module are fine-tuned.
Training Objective
The training objective of VolCano is unified as a causal modeling loss:
$\mathcal{L(\theta)}=-\mathbb{E}_{x_{i}\sim D}\Big{[}\frac{1}{|v_{i}|}\sum_{t\in v_{i}}\log P_{\theta}(x_{t}|\{x_{j}\}_{j<t})\Big{]}$ (1)
where $x_{i}$ is the constructed sequence in $D$, the training dataset. $P_{\theta}$ is the probability distribution modeled by the model and $\theta$ denotes the parameters. For each sequence $x_{i}$, $v_{i}$ indicates the set of indexes for tokens that need to be predicted. In practice, $v_{i}$ only contains textual tokens, including special tokens, coordinates, and natural text. In terms of pre-training data, all textual tokens are included in $v_{i}$, whereas for instruction-tuning data, only tokens in the responses are considered in $v_{i}$.
4 Experiments
4.1 Experiment Settings
Implementation Details
In this paper, we build VolCano with the pre-train ViT-L/14 CLIP [40] visual encoder and Mistral-7B [16] as the language backbone. Compared to Vicuna [8] and LLaMA [45], Mistral offers superior reasoning abilities in text, making it a better choice for modeling VoCoT. The input resolution of images are set to $336\times 336$. Experiments are conducted using 8 NVIDIA A100 GPUs. The model is trained for one epoch in each stage, the complete training process takes approximately 90 hours. Detailed parameter settings are provided in Appendix A.3.
Tasks and Datasets
To validate the effectiveness and versatility of the multiple-reasoning paradigm formulated by VoCoT, we adopt different tasks across various scenarios for assessment:
  • General VQA comprehensively evaluates various capabilities of LMMs by integrating different types of instructions and questions. We consider the commonly adopted GQA [15], MMBench [33], and SEED [21] benchmarks. For SEED, we observe that different evaluation methods are used by different models. To conduct a fair comparison, we re-evaluate the compared models with the likelihood evaluation method described in the original paper [21].
  • Visual spatial reasoning requires models to distinguish the spatial relationship between objects. VSR [28] and EmbSpatial [10] are utilized for assessment.
  • Hallucination tasks. Since hallucination significantly hinders effective reasoning. We utilize POPE [24] and AMBER [48] to diagnose the reliability of the generated outputs.
  • Compositional tasks in multi-modal problems require solution logic that includes multi-step analysis and reasoning. CLEVR [17], V-Star [53], and Winoground [44] are applied. Following Reform-Eval [25], we formulate Winoground as a binary classification task of selecting the correct description from two similar captions based on an image.
Except for AMBER, where CHAIR [48] is the evaluation metric, accuracy is reported for other datasets. For details on the evaluation datasets and evaluation methods, please refer to Appendix A.4.
Baseline Models
We compare VolCano with existing LMMs based on single-step reasoning with ~7B parameters, including BLIP-2 [22], InstructBLIP [9], Shikra [6], mPLUG-Owl2 [59], MiniGPT-v2 [5], Qwen-VL-Chat [2], and LLaVA-1.5 [29]. Furthermore, based on our framework and data with VoCoT-Instruct-80K removed, we construct a single-step reasoning baseline model, VolCano-SE, for a strict comparison. The models listed before, including VolCano, take a single image as input. Additionally, we include another series of SOTA LMMs that enhance input resolution by splitting a single image into multiple sub-images: LLaVA-1.6 [30], Deepseek-VL [34], Monkey [27].
Notations
[…]
4.2 Main Results
Table 2: Comparison with SOTA LMMs on 10 benchmarks.
Model General VQA Spatial Reasoning Compositional Tasks Hallucination
[…]
Models with single-image inputs
BLIP-2-7B $224^{2}$ 1.3B 41.00 - - 50.90 29.50 - 37.70 - - -
InstructBLIP-7B $224^{2}$ 1.3B 49.20 36.00 - 52.10 29.00 - 34.02 - 72.10 8.80
Shikra-7B $224^{2}$ 0.3B - 58.80 - - 32.60 - - - 83.10 -
mPLUG-Owl2-7B $448^{2}$ 0.3B 56.10 64.50 59.99 - 36.72 43.22 36.12 63.38 - 10.60
MiniGPT-v2-7B $448^{2}$ 1.3B 60.10 55.14 51.50 62.90 45.10 46.23 33.19 62.00 80.50 -
Qwen-VL-Chat $448^{2}$ 1.9B 57.50 60.60 64.70 - 38.68 53.20 45.80 - 84.70 5.50
LLaVA1.5-7B $336^{2}$ 0.3B 62.00 64.30 53.80 64.24 42.43 43.73 48.68 55.31 84.50 7.80
VolCano-SE $336^{2}$ 0.3B 59.91 61.15 54.15 63.42 36.14 51.70 44.96 64.00 84.50 6.70
VolCano $336^{2}$ 0.3B 64.40 68.10 64.50 67.18 58.29 56.17 58.40 68.37 86.50 4.60
Models with multiple-image inputs
LLaVA1.6-7B $672^{2}$ 0.3B 64.20 68.40 66.15 66.86 56.82 50.35 58.80 64.88 86.90 -
LLaVA1.6-7Bm $672^{2}$ 0.3B 64.80 69.00 67.72 63.77 56.55 51.85 60.08 65.75 86.70 -
Deepseek-VL-7B $1024^{2}$ 0.4B - 71.32 70.40 67.51 41.77 48.77 62.18 64.88 85.77 -
Monkey $896^{2}$ 1.9B 60.70 61.95 67.58 62.93 32.91 46.33 67.23 68.63 82.57 -
GPT-4V $2048^{2}$ 75.80 71.60 68.24 26.65 51.90 55.00 83.75 82.00 4.60
Table 2 presents a thorough evaluation of existing LMMs. Several insights can be gleaned: (1) Among models with single-image inputs, VolCano achieves the best performance in most datasets except SEED-Image, with the advantage being more pronounced in complex tasks. (2) Introducing VoCoT in LMMs effectively alleviate the hallucination issue. Section 4.3 delves deeper into how VoCoT contributes to reliable and visually grounded reasoning. (3) Comparing VolCano with LMMs that enhance input resolution by introducing multiple-image inputs, we observe that these methods primarily improve the performance in general VQA and V-Star, as V-Star provides high-resolution input images [53]. However, in tasks that require complex reasoning, the improvement brought by higher resolutions becomes less significant. VolCano either exceeds or perform comparably with these models in such tasks, indicating the superiority of introducing multi-step reasoning over enriching the input information in these scenarios. We leave exploring combining these two vertical research directions—enhancing input resolution and introducing multi-step reasoning–as future work. Overall, the experimental results validate the effectiveness of VoCoT-based multi-step reasoning in various scenarios, particularly in tasks demanding complex reasoning.
4.3 Comparison between CoTs in Different Formats
In this section, we demonstrate the effectiveness of the VoCoT format by comparing it with different CoT formats: (1) Zero-Shot CoT directly prompts VolCano-SE model to include a object-centric reasoning path without training; whereas (2) Text CoT represents objects with only text descriptions; (3) Coor. CoT augments Text CoT with coordinates; and (4) Sub-Img CoT encodes sub-images as visual representations of objects rather than using RefBind, during the training process.
Table 3 lists the results. Firstly, the zero-shot multi-step reasoning capability of VolCano-SE is limited and likely to exhibit hallucinations, highlighting the necessity to construct visual CoT tuning data. Secondly, CoT expressed only in texts is also affected by hallucinations, handling spatial reasoning well where each type of object appears only once, but failing to manage more complex scenarios. Thirdly, introducing coordinates grounds the thoughts to visual signals, mitigating hallucination and improving performance across various tasks. Furthermore, the representations obtained by RefBind effectively help the model utilize visual signals of objects. However, simply encoding sub-images may backfire by including redundant information. Overall, VoCoT is the most suitable format.
Table 3: Comparison between different CoT formats. The “Obj-Format” column indicates the representation format of objects. T, C, S, and R are short for texts, coordinates, sub-images, RefBind.
[…]
Zero-Shot CoT < T > 56.79 52.47 51.70 45.32 57.20 65.00 67.50 52.20 6.70
Text CoT < T > 63.36 59.20 49.60 47.90 68.49 65.75 84.63 49.30 5.50
Coor. CoT < T, C > 64.32 58.59 54.42 53.78 66.86 65.87 85.47 47.80 4.30
Sub-Img CoT < T, C, S > 61.29 54.10 51.85 63.45 66.23 58.87 85.77 47.80 4.60
VoCoT < T, C, R > 64.50 58.29 56.17 58.40 67.18 68.37 86.50 51.00 4.60
4.4 Ablation on Constructed Datasets
In Table 4, we explore the role of three types of data in VoCoT-Instruct-80K. The results implies that: (1) Type 1, the GQA-based data, is precise but limited in terms of diversity. Models trained solely on Type 1 data produces the fewest hallucinations but struggle to handle diverse questions. (2) Type 2 and Type 3 data effectively help the model generalize across various instructions. Nevertheless, removing Type 1 data from the training set increases the risk of hallucinations. (3) Introducing multi-modal interleaved data in Stage 2 significantly improve the performance. In summary, three types of data should be jointly utilized, and interleaved data need to be incorporated for pre-training.
Table 4: Ablation of VoCoT-formatted data on stage 2 and stage 3 training process.
[…]
Type 1 Type 2 Type 3
63.63 47.78 56.07 57.14 68.90 66.00 84.80 33.70 1.80
65.45 58.21 54.93 61.34 66.85 65.00 80.16 48.90 5.00
64.20 57.24 54.17 53.36 68.98 66.87 86.50 48.90 4.40
64.50 58.29 56.17 58.40 67.18 68.37 86.50 51.00 4.60
61.62 57.22 49.18 53.36 67.13 63.62 85.60 46.90 5.50
4.5 Further Analysis
VoCoT enhances performance in complex questions
Figure 4: Relationship between performance and the number of reasoning steps required by the questions.
Figure 4 compares the performance of VolCano-SE and VolCano on questions of varying difficulty in CLEVR [32]. Based on the fitted curves and confidence intervals, it is evident that as the number of required reasoning steps increases, the advantage of multi-step reasoning becomes more pronounced.
Grounding Capabilities of VolCano
Table 5: Referring expression results.
Model RefCOCOg CLEVR Ref
Method Res. val test
Shikra-7B $224^{2}$ 82.3 82.2 -
MiniGPT-v2-7B $448^{2}$ 84.2 84.3 24.9
Qwen-VL-Chat-7B $448^{2}$ 85.9 86.3 22.4
LLaVA1.5-7B $336^{2}$ 70.5 71.0 6.7
VolCano-7B $336^{2}$ 83.1 84.4 34.0
In this section, we attempt to separately analyze two aspects of capabilities in VolCano: visual grounding and reasoning. Referring expression task in RefCOCOg [18] and CLEVR-Ref [32] is adopted to evaluate the former capability. According to Table 5, VolCano with lower input resolution demonstrates comparative performance on RefCOCOg which is an in-domain dataset. In the out-domain and complex dataset, CLEVR-Ref, VolCano significantly outperforms other models. These results indicate that VolCano possesses strong and generalizable grounding ability.
Reasoning Capabilities of VolCano
Regarding the reasoning process, we divide it into two sub-processes: analysis and judgment, where the former constructs reasoning paths and the latter provides conclusions. We conduct experiments on VSR: VolCano serves as the analyzer to provide reasoning paths based on which the conclusions are drawn by different judgers. In VSR, there is one-to-one correspondence between objects in the image and textual descriptions, LLMs can effectively make judgments based on the textual reasoning paths generated by VolCano after removing the coordinates and visual information. In addition, we construct VolCanoV, a Vicuna-based variant of VolCano, to investigate the impact of language backbones. According to Table 6, a more powerful judger can more effectively utilize the reasoning paths generated by Vocalno. When using VolCano as the analyzer and GPT-4 as the judger, the performance even surpasses that of GPT-4V. This indicates that VoCalno possesses the capability to provide effective reasoning paths, but its judgment ability is limited. Furthermore, it can be observed that the judgment ability of VolCano is mainly limited by the backbone. Generally, the experiments highlight the importance of the language backbone and reveal the potential of further applying VoCoT on stronger LLM backbones.
Table 6: Performance with different analyzers and judgers on VSR dataset.
Analyzer VolCanoV VolCano GPT-4V
Judger VolCanoV Vicuna-1.5 Mistral VolCano Vicuna-1.5 Mistral GPT-4
Accuracy 63.5 64.5 67.2 67.2 65.1 67.8 73.8 68.24
Case Study
As shown in Figure 5, in AMBER, VolCano provides a visually grounded description with no hallucinations. In CLEVR, VoCalno can infer effective multi-step reasoning paths towards the answer. Please refer to Appendix B.1 for more cases from different datasets.
5 Related Works
5.1 Large Multi-Modal Models
Architecture of Large Multi-Modal Models
A vast amount of research has emerged, focusing on adapting large language models (LLMs) to handle multi-modal tasks. Initially, researchers treated large language models (LLMs) as intelligent agents capable of using various tools. They trained these models to invoke fundamental vision models, enabling them to complete multi-modal tasks such as object detection, image description, and visual question answering [52; 43; 56]. More recent works have focused on directly aligning the visual and textual modalities. To achieve this, researchers design various connection modules, including linear projections [31; 11; 5; 30; 29; 34; 6; 26], Q-Former [22; 14; 62; 9], and cross-attention-based re-sampler (abstractor) [1; 58; 2; 27; 59].
Visual Instruction Data Construction
To enhance the instruction-following abilities of large language models (LLMs) in multi-modal contexts, a significant line of research has focused on converting existing task-oriented datasets into instructional data [9; 29; 23]. Utilizing models akin to GPT-4, such as MiniGPT4 [62], researchers bootstrap high-quality data from the model outputs, and LLaVA [30] transforms image-caption pairs into dialogue data. Further efforts are directed towards improving both the quality and scale of the generated data [4; 61; 47].
5.2 Multi-Step Reasoning within Large Pre-trained Models
CoT in Large Language Models
Chain of Thoughts (CoT) is a series of prompting techniques designed to facilitate large language models (LLMs) in addressing complex problems by guiding them through intermediate steps. CoT is first proposed through in-context learning [51], followed by the introduction of zero-shot CoT [19], Auto-CoT [60] and self-consistency mechanism [50]. Subsequently, CoT are adapted to consider complex formats [57; 3; 7].
Visually Enhanced Reasoning
To address complex multi-modal problems, Shikra [6] initially explores the potential of applying CoT to specific tasks with LMMs. SoM [54] and Scaffolding [20] incorporate additional structural information into images, such as segmentation maps and dot grids. These enhancements help the model utilize these anchors for reasoning. However, these approaches are specifically designed for GPT-4V, which has strong visual perception capabilities. V-Star [53] designs a reasoning logic for visual search in high-resolution images, which can be considered a simple two-step CoT. Overall, CoT has not been comprehensively explored in LMMs and there lack appropriate reasoning formats that could be generalized to various scenarios and tasks.
Figure 5: Qualitative analysis with cases from AMBER and CLEVR. Hallucinations are highlighted.
6 Conclusion
In this paper, we introduce VoCoT, a visually-grounded and object-centric Chain of Thoughts (CoT) format, designed to assist large multimodal models (LMMs) in multi-step reasoning. Furthermore, we propose a pipeline to construct VoCoT-formatted data from existing resources, leading to the creation of VoCoT-Instruct-80K. Based on this data, we develop VolCano, a model designed to perform multi-step reasoning effectively and efficiently using the VoCoT format. Comprehensive experiments demonstrate the effectiveness of our approach.
Appendix A Supplementary Details
A.1 Data Construction Details
Construction Methods
As for the Type 1 data construction, we utilize a rule-based conversion method, the mapping rules used are listed in Table 11. In terms of Type 2 and Type 3 data, the prompts for GPT-4V are respectively shown in Table 12 and  13. We use in-context learning to let GPT-4V generate thought with multi-step reasoning path in VoCoT format.
Quality Control
For Type 1 data generated by rule based mapping, the quality is controled by the original source, namely GQA. We do not perform additional quality control. For Type 2 and Type 3 data generated by GPT-4V, we manually check 200 samples, we find if GPT-4V think the information we provide is not enough to generate correct reasoning path, model response will contains error messages. All error message have patterns in common which may contains following phrase: “From the object information provided", “provided object information", “From the bounding boxes provided", “provided bounding boxes", “provided object bounding boxes", “object bounding boxes provided", “object bounding box information provided", “bounding box information provided", “object information provided", “given object information", “the bounding boxes provided", “From the object bounding box information", “provided object bounding box information". We remove the sample containing these patterns, leaving 8k samples after the filtering.
A.2 Training Data Details
Table 7: Data Mixture in 3 Stages of VolCano.
Stages Data Type Source Size
Stage 1 Image-Caption LLaVA-Pretrain [31; 29] 558k
Stage 2 Image-Caption ALLaVA-Caption [4] 695k
Grounded Image Caption GRIT [38] 756k
Flickr30k-Entities [39] 148k
Multimodal Document MMC4 [63] 890k
Stage 3 Visual Instruction LLaVA [29] 612k
Referring Expression Shikra-RD [6] 6k
RefCOCO [18] 42k
RefCOCO+ [18] 42k
RefCOCOg [18] 42k
g-RefCOCO [13] 79k
VoCoT This Work 80k
We present our data mixture details in Table 7. In Stage 1, we use LLaVA-pretrain [29] dataset for projector alignment which contains 558k image caption pairs. In Stage 2, to better adapt to multi-modal interleaved sequences and visually grounded object representations with Refbind, we mix three types of data: (1) multimodal documents, (2) grounded image captions, and (3) high-quality image captions. Multimodal documents data is sample from MMC4 [63] by choosing which average similarity score between image and sentence before each image is larger than 0.3. Each multimodal document have multiple image, we remove samples with more than 6 images. Grounded image captions is from GRIT [38] and Flickr30K Entities [39]. For GRIT, we filter samples with clip score is larger than 0.35. We also use high quality image caption from ALLaVA [4] which is generated by GPT-4V and provide details description. In Stage 3, we remove samples from LLaVA-Instruct that meet two criteria: sourced from RefCOCO, and sourced from VG where the object sub-image size is less 50, the reason is that we find extremely small regions in VG are probably with low quality.
A.3 Model & Training Details
We supplement the hyper-parameters in each stage in Table 8.
Table 8: The detailed training hyper-parameters of VolCano
Configuration Alignetent Multi-modal Interleaved Instruction Tuning
Visual Encoder OpenAI-CLIP ViT-L/14 OpenAI-CLIP ViT-L/14 OpenAI-CLIP ViT-L/14
Backbone Init Mistral-Chat-v0.2-7B Stage1 Stage2
Optimizer AdamW AdamW AdamW
Optimizer Hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e^{-6}$ $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e^{-6}$ $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e^{-6}$
Global batch size 256 128 128
Peak learning rate of LLM 5e-5 1e-5 1e-5
Learning rate schedule Cosine Cosine Cosine
Training Epochs 1 1 1
Warm-up ratio 0.03 0 0
Weight decay 0.0 0.0 0.0
Gradient clipping 1.0 1.0 1.0
Input image resolution 336 * 336 336 * 336 336 * 336
Input sequence to LLM 2048 3072 3072
Numerical precision bfloat16 bfloat16 bfloat16
GPU Usage 8 NVIDIA A100 8 NVIDIA A100 8 NVIDIA A100
Training Time 12h 48h 30h
Following Kosmos2 [38], we introduce a special token “<grounding>” at the beginning of the sequence to control whether to require the model to produce visually grounded description.
Also notice that VolCano is able to perform both single-step and multi-step reasoning. Following [19], we introduce a prompt “Answer the question and include the reasoning proess. Locate key objects and provide bounding boxes in your thoughts.” as a trigger to tell the model whether to use VoCoT or not during the generation process.
A.4 Evaluation Details
In this section, we introduce the details in the evaluation procedure.
Table 9: Information of Evaluation Benchmarks.
Category General VQA Spatial Reasoning Hallucination
Datasets GQA MMBench SEED VSR EmbSpat. POPE AMBER
Split testdev_balanced DEV Image test unseen test adversarial generative
Data Amount 12578 4329 14233 1222 3625 3000 1004
Table 10: Supplementation of Table 9.
Category Compositional Tasks Referring Expression
Datasets V-Star Wino CLEVR RefCOCOg CLEVR-Ref
Split - test val val (umd) test (umd) -
Data Amount 238 800 6000 4896 9602 2000
A.4.1 Benchmark Details
General VQA Benchmarks
In GQA, we utilize the “testdev_balanced” split following [31]. As for MMBench, we adopt the “DEV” split for the evaluation efficiency. In terms of SEED, we only consider the subset that the visual inputs are images. For GQA, we append a prompt “Please answer in a word or short phrase” to require models produce concise outputs.
Spatial Reasoning Benchmarks
For VSR, we utilize the unseen test split for zero-shot evaluation. For each sample, a description is provided and the model is required to distinguish if the claims is supported by the image. We use the prompt “Is there a event {description} in the image?” for this dataset. With respect to EmbSpatial, we use the test split for assessment.
Hallucination Benchmarks
For POPE, we consider the adversarial subset since it is the most challenging split. In AMBER, we leverage the generative task which asks the model to describe the image. All prompts are directly adopted from the original datasets.
Benchmarks for Compositional Tasks
For CLEVR, we utilize the val split. Because the original CLEVR validation set is too large, we categorize the data the into six types based on the question type: count, yes/no, shape, material, size, and color. We sample 1k questions from each category as test samples and construct a multiple-choice candidate set based on the feasible answers in the dataset. We will also open-source this subset. For Wino, we utilize the test set and consider it as a caption selection multiple-choice question, the prompt is designed as “Please describe the image.”. As for V-Star, we directly use the V-Star benchmark.
Referring Expression Benchmarks
For RefCOCOg, we utilize the val and test splits which is provided by umd. As for CLEVR-Ref, we use the provided set for evaluation. We design a prompt as "Can you locate phrase in the image?" where “{phrase}” is the target query.
Please see Table 10 and Table 9 for the dataset splits and scales.
A.4.2 Evaluation Methods
All evaluation benchmarks adopted in this paper can be divided into three categories based on the task formulation: multiple-choice questions, open-ended generation, referring expression.
Multiple-Choice Question
For multiple-choice questions, we utilize the likelihood-based evaluation method, which is also known as the perplexity-based method. These methods are widely adopted in evaluating LMMs [21; 22; 9; 25]. The key idea is to select the option with the highest generated likelihood, please refer to these papers for the detail. If VoCoT is utilized, the likelihood is computed based on the question, image, and the genrated reasoning path.
Open-Ended Generation
For GQA, we use the evaluation script provided by LLaVA [29] for a fair comparison. As for VSR and POPE, we require the model to answer in yes and no, enabling us to evaluate the correctness with exact match. With respoect to AMBER, we use the official evaluation method to assess the hallucinations in the generated descriptions.
Referring Expression
We first extract the predicted boxes from the outputs based on rules, then IoU between the ground truth box and predicted box is calculated. If the IoU is larger than 0.5, it is considered as a correct prediction [18].
Further Analysis Setup
In the reasoning capability assessment part in Section 4.5, to leverage LLM as the judger model. We utilize a prompt “There is a image, {reasoning path}, please determine whether {description}, please answer yes or no.”, where the reasoning path are generated by the analyzer (with the coordinates and visual information removed), descriptions are the target description in VSR. If the model chooses not to predict, we consider the prediction as “no”.
Appendix B Supplementary Analysis and Discussion
B.1 Case Study
Different from the black-box single-step reasoning paradigm, VolCano produces interpretable responses with the reasoning paths in text. Please see Figure 67891011 for cases from representative datasets.
B.2 Limitation and Future Work
Although VoCoT can be applied to assist LMM in various tasks, there still exists limitations:
  • VoCalno is limited in terms of understanding certain visual-related concepts, such as spatial relationship and textures. We attribute this to insufficient training on vision-related data. In the future, richer data and text descriptions related to these concepts should be introduced in training LMMs.
  • Currently, VoCoT is not applicable to multi-image inputs. This issue could be addressed by including additional special tokens to enable VoCoT distinguish between objects in different images. We leave it as the future work.
  • Limited by the cost of data construction, VoCoT-Instruct can not be effectively scaled up currently. We hope to explore more efficient VoCoT data construction methods in the future.
  • VolCano is currently limited with respect to the model scale and input resolution. However, the experimental results reveal the potential of applying VoCoT in larger backbones, which is the future direction for the development of VolCano.
Table 11: Mapping rule for transferring SQL-like query statement to string in GQA Source Type Data construction.
Opeation Mapping Rule
relate: sub, relation, obj Check the {subject} that is {arg2} {object}.
same: attribute [obj1,obj2] The question ask if the two objects has same {attribute}.
Check if they have same {attribute}.
common: [obj1,obj2] The question ask the common attribute of the two objects.
different: attribute, [obj1, obj2] The question ask if the two objects has different {attribute}
and: [obj1, obj2] The question ask about ’and’ relation.
select: obj1 Find {obj1}
exist: ? obj1 It doesn’t exist. if obj1 is not in annotation else It exist
verify: attribute,value, obj1 Verify if the {attribute} of {obj1} is {value}.
or: [obj1, obj2] The question ask about ’or’ relation.
choose: obj1, attribute, value1, value2, obj2 Think {obj1}’s {attribute} is {value1} or {value2} of {obj2}.
choose: obj1, attribute, value1, value2, Think {obj1}’s {attribute} is {value1} or {value2}.
Table 12: Prompt and one in-context sample for VQA-Based Source Type Data generation.
messages = [ {"role":"system", "content": f"""You are an excellent generator of image QA reasoning processes based on question-answer pairs and object information represented by object bounding boxes ($x_{left_{t}op}$, $y_{left_{t}op}$, $x_{right_{d}own}$, $y_{right_{d}own}$).Your task is to generate reasoning process based on the questions and answers you are given. The reasoning process should include the reasoning path, relevant object bounding boxes, and inference clues, including but not limited to the object’s number, location, and your own background knowledge. The object in your reasoning path must annotate with object bounding box. The bounding box must come from the object information given by the user, please do not detect it yourself! ! ! ! Don’t mention object information directly, just annotate it with bounding boxes.When you refer to the information in prompt, the text should show that you did not know the answer in advance, but that you reasoned it out yourself. And don’t directly say that something doesn’t appear in the information provided. Don’t mention anything in the prompt in your reply, and don’t mention bounding boxes in the generated reasoning process.You will follow instructions to the best of your ability. Your response should follow the following format: "Thought":"""""}, [Uncaptioned image] {"role": "user","content": [{"type": "text","text": "[IMAGE1]:[Object Info]: Coffee1: [0.04, 0.25, 0.20, 0.28], Bean: [0.22 , 0.12 , 0.29 , 0.32 ], Vegetable: [0.52, 0.11, 0.29, 0.31 ], Coffee2: [0.76, 0.25, 0.21, 0.27 ], Yam: [0.69, 0.48, 0.27, 0.34 ], Burrito: [0.34, 0.40, 0.35, 0.44 ], Orange: [0.05, 0.50, 0.27, 0.34 ], Chopsticks: [0.61, 0.69, 0.22, 0.31] [question]: What is the food with the most Vitamin C in this image?" [answer]: Orange."},{"type": "image_url", "imag_url": {"url": f"data:image/jpeg;base64,{image1}"}}]}, {"role": "assistant", "content":”’"thought": "From the picture, you can see two cups of coffee [0.04, 0.25, 0.20, 0.28],[0.76, 0.25, 0.21, 0.27 ], one portion of bean[0.22 , 0.12 , 0.29 , 0.32 ], one portion of vegetables[0.34, 0.40, 0.35, 0.44 ], one portion of yam[0.69, 0.48, 0.27, 0.34 ], one burrito[0.34, 0.40, 0.35, 0.44 ], and a plate of oranges[0.05, 0.50, 0.27, 0.34 ]. Among these foods, bean[0.22 , 0.12 , 0.29 , 0.32 ] contains protein, yams[0.05, 0.50, 0.27, 0.34 ] and burrito[0.34, 0.40, 0.35, 0.44 ] are rich in starch, vegetables[0.34, 0.40, 0.35, 0.44 ], and oranges[0.05, 0.50, 0.27, 0.34 ] are foods that may contain vitamin C, but oranges[0.05, 0.50, 0.27, 0.34 ] have a higher vitamin C content, so oranges[0.05, 0.50, 0.27, 0.34 ] are foods that contain more vitamin C."”’},]
Table 13: Prompt and one in-context sample for Image-Only Source Type Data generation.
messages = [ {"role":"system", "content": f"""You are an excellent image describer and question-answer generator based on the image and object information which is represented by object bounding box ($x_{left_{t}op}$, $y_{left_{t}op}$, $x_{right_{d}own}$ , $y_{right_{d}own}$). You have three tasks in total. Your first task is to ask a complex question that requires close inspection of the image and strong reasoning ability to answer. Your second task is to answer the question you raised solely based on the given image. Your third task is to generate the reasoning thought. The reasoning thought should contain the reasoning path, relative object bounding box and inference clue, include but are not limited to the object numbers, location and background knowledge from yourself. The bounding box must come from the user given object information, Do not detect by yourself !!!! Do not mention the object infomation directly, just annotate with bounding box. When you ask questions, try to find the most valuable information in the picture to ask about, and ask a question that is relevant to that information. When you ask questions, do not involve violence, advertisement, possible invasion of privacy, or questions that may cause discomfort. Do not mention anything from the prompt in your response and Do not mention bounding box in your generated question. You will follow the instructions to the best of your ability. Your response should follow the following format: "question":"","answer":"","Thought":"""""}, [Uncaptioned image] {"role": "user","content": [{"type": "text","text": "[IMAGE1]:[Object Info]: Coffee1: [0.04, 0.25, 0.20, 0.28], Bean: [0.22 , 0.12 , 0.29 , 0.32 ], Vegetable: [0.52, 0.11, 0.29, 0.31 ], Coffee2: [0.76, 0.25, 0.21, 0.27 ], Yam: [0.69, 0.48, 0.27, 0.34 ], Burrito: [0.34, 0.40, 0.35, 0.44 ], Orange: [0.05, 0.50, 0.27, 0.34 ], Chopsticks: [0.61, 0.69, 0.22, 0.31]"},{"type": "image_url", "imag_url": {"url": f"data:image/jpeg;base64,{image1}"}}]}, {"role": "assistant", "content":”’"question":"What is the food with the most Vitamin C in this image?", "answer": "Orange.", "thought": "From the picture, you can see two cups of coffee [0.04, 0.25, 0.20, 0.28],[0.76, 0.25, 0.21, 0.27 ], one portion of bean[0.22 , 0.12 , 0.29 , 0.32 ], one portion of vegetables[0.34, 0.40, 0.35, 0.44 ], one portion of yam[0.69, 0.48, 0.27, 0.34 ], one burrito[0.34, 0.40, 0.35, 0.44 ], and a plate of oranges[0.05, 0.50, 0.27, 0.34 ]. Among these foods, bean[0.22 , 0.12 , 0.29 , 0.32 ] contains protein, yams[0.05, 0.50, 0.27, 0.34 ] and burrito[0.34, 0.40, 0.35, 0.44 ] are rich in starch, vegetables[0.34, 0.40, 0.35, 0.44 ], and oranges[0.05, 0.50, 0.27, 0.34 ] are foods that may contain vitamin C, but oranges[0.05, 0.50, 0.27, 0.34 ] have a higher vitamin C content, so oranges[0.05, 0.50, 0.27, 0.34 ] are foods that contain more vitamin C."”’},]
Figure 6: Cases on CLEVR dataset.
Figure 7: Cases on GQA dataset.
Figure 8: Cases on Seed-Bench dataset. Seed-Bench provide a question and a list of answer options for each sample.
Figure 9: More Cases on Seed-Bench dataset.
Figure 10: Cases on VSR dataset. VSR dataset provides a statement for each image. The task is to judge whether the statement is right about this image.
Figure 11: Cases from the AMBER dataset. The task is to describe each image. We both present our VolCano and VolCano-SE responses for these cases. VolCano-SE is trained without VoCoT data. The underline phrase is hallucination generated by VolCano-SE.