### v3

# VoCoT: Unleashing Visually Grounded Multi-Step Reasoning in Large Multi-Modal Models

Abstract
While large multi-modal models (LMMs) have exhibited impressive capabilities across diverse tasks, their effectiveness in handling complex tasks has been limited by the prevailing single-step reasoning paradigm. To this end, this paper proposes VoCoT, a multi-step Visually-grounded object-centric Chain-of-Thought reasoning framework tailored for inference with LMMs. VoCoT is characterized by two key features: (1) object-centric reasoning paths that revolve around cross-modal shared object-level information, and (2) visually grounded representation of object concepts in a multi-modal interleaved and aligned manner, which effectively bridges the modality gap within LMMs during long-term generation. To adapt LMMs in reasoning with VoCoT, we further construct an instruction-tuning dataset. By combining VoCoT with the prevalent open-source LMM architectures, we develop a VoCoT-based model, VolCano. With only 7B parameters and limited input image resolution, VolCano demonstrates excellent performance across various scenarios. In benchmarks like CLEVR and EmbSpatial, which highly require complex reasoning capabilities, VolCano outperforms SOTA models, including powerful GPT-4V. Related code, models, and datasets are released in https://github.com/RupertLuo/VoCoT.
VoCoT: Unleashing Visually Grounded Multi-Step Reasoning
in Large Multi-Modal Models
Zejun Li1222Equal contribution., Ruipu Luo2222Equal contribution., Jiwen Zhang1, Minghui Qiu2, Xuanjing Huang3, Zhongyu Wei1,4,5111Corresponding authors. 1School of Data Science, Fudan University 2ByteDance 3School of Computer Science, Fudan University 4Research Institute of Intelligent and Complex Systems, Fudan University 5Shanghai Innovation Institute {zejunli20,zywei}@fudan.edu.cn
1 Introduction
In recent years, the success of large language models (LLMs) OpenAI (2023a, b) has been gradually extended to the multi-modal domain. By equipping LLM backbones (Touvron et al., 2023a, b; Chiang et al., 2023) with visual encoders (Radford et al., 2021) and efficient cross-modal alignment through generative training on image-text data (Liu et al., 2024b; Schuhmann et al., 2021), the constructed large multi-modal models (LMMs) possess the capabilities to perceive visual signals and engage in dialogue with users in multi-modal contexts Liu et al. (2024b); Dai et al. (2023); Bai et al. (2023).
Despite the potential demonstrated by LMMs to serve as unified and versatile foundations, even models like GPT-4V struggle in composite tasks requiring complex analysis Yang et al. (2023b); Wu and Xie (2023), such as spatial reasoning Du et al. (2024). We attribute this phenomenon to a major limitation of current LMMs: the prevailing single-step question-to-answer (Q2A) inference paradigm that directly generates answers based on questions Dai et al. (2023); Liu et al. (2023b). As illustrated in Figure 1, correctly answering the question relies on analyzing the actions and relationships of multiple objects and thinking step-by-step, which is almost impossible to accomplish in a single-step prediction. Moreover, the single-step Q2A paradigm obscures the problem-solving process, limiting the interpretability of the LMM outputs. Conversely, in the language domain, the chain-of-thought (CoT) paradigm, which involves multi-step reasoning, has been widely explored in LLMs Kojima et al. (2022); Wei et al. (2022), indicating a promising way for enhancing LMMs.
Figure 1: An example to compare different inference paradigms in LMMs. (a) A visual question that requires complex reasoning. (b) The conceptual object-centric reasoning path constructed to solve the problem. (c) Outputs of GPT-4V and the proposed VolCano. Hallucination is included in the output of GPT-4V. VoCalno performs multi-step reasoning in the VoCoT format. In the reasoning path, key objects are highlighted and colors indicate the correspondence between object descriptions and the grounded regions in the image. “[box]” represents the coordinates of mentioned objects. Visual representations of objects are omitted for brevity.
However, for complex contexts where multi-modal information coexists, constructing effective multi-step reasoning paths faces several challenges: (1) Difficulty in integrating reasoning anchors within multi-modal contexts. Textual CoTs mainly extract key information from contexts, such as entities, as anchors and conduct multi-step reasoning around these anchors Yao et al. (2024). In multi-modal contexts, the anchor information is further required to be concepts shared between images and texts and establish connections between modalities. Existing works either supplement the image with additional information (such as segmentation maps Yang et al. (2023a) and dot grids Lei et al. (2024)) as anchors, but such information can only be effectively utilized by GPT-4V, or they roughly consider a single region as the anchor through a search-based approach Shao et al. (2024), failing to model complex multi-object interactions. (2) Limited grounding capabilities of LMMs. During the generation process, LMMs may fail to ground textual descriptions to the corresponding visual information, resulting in erroneous information generated. For example, GPT-4V incorrectly ground the target person to the waiter in Figure 1. The risk of hallucination Li et al. (2023d); Wang et al. (2023b) further hinders effective multi-step reasoning.
To address these challenges, we introduce a framework to empower LMMs for effective and reliable multi-step reasoning. We propose VoCoT, Visually grounded object-centric Chain of Thought. VoCoT is a CoT format that is compatible with LMM inference: (1) As illustrated in Figure 1 (a, b), objects serve as fundamental semantic units in both images and text, effectively bridging multi-modal information. Therefore, VoCoT leverages objects as anchors for reasoning. LMMs are encouraged to conduct multi-step analysis on the properties of key objects, as well as the relationships between them, ultimately reaching a conclusion. (2) To ensure the reliability of reasoning paths, VoCoT represents objects in a visually grounded format: a tuple of <textual description, coordinates, corresponding visual representations>. Models are required to explicitly ground objects in images by generating coordinates for them. Visual representations of objects are supplemented to enhance the cross-modal relevance in reasoning paths. This design mimics the habit of human, where we continuously reference the visual information of an object in the image when we mention it. (3) We propose a RefBind mechanism to efficiently obtain the representations of objects without extra computation. Specifically, RefBind indexes the representation of each object from the image representation based on its coordinates. Generally, VoCoT constructs multi-modal interleaved reasoning paths where cross-modal aligned anchors are incorporated as shown in Figure 1 (c).
Nevertheless, there is a significant disparity between VoCoT and the formats of existing visual instruction data. To this end, we further construct a dataset, VoCoT-Instruct-80K, to train LMMs for reasoning in the format of VoCoT. VoCoT-Instruct-80K is built on multiple data sources: (1) Verbalizing structured reasoning paths from GQA Hudson and Manning (2019). (2) Supplementing visual QA pairs with thought processes. (3) Constructing complex questions and reasoning paths from images annotated with objects. By curating a wide range of data and leveraging assistance of GPT-4V, the presented dataset maintains both diversity and consistency in the desired format.
Based on the introduced VoCoT framework and dataset, we develop VolCano, a Visually-grounded multi-modal Chain-of-thought reasoning model. With only 7B parameters and $336^{2}$ input resolution, VolCano excels in various scenarios and even surpasses GPT-4V on benchmarks like CLEVR Johnson et al. (2017) and EmbSpatial Du et al. (2024) that highly require complex reasoning.
2 Visually-grounded Object-centric CoT
In this section, we explain how to enable LMM to perform multi-step reasoning in the format of visually-grounded object-centric chain-of-thought (VoCoT). In Section 2.1, we elaborate on the formulation of VoCoT. In Section 2.2, we present how to transform existing data resources into instruction-tuning datasets aligned with the VoCoT format.
2.1 VoCoT Formulation
VoCoT requires LMMs to perform step-by-step reasoning based on the provided context. Following textual CoTs Kojima et al. (2022); Wei et al. (2022); Wang et al. (2022), the reasoning logic in VoCoT is primarily expressed in text but is not limited to specific formats. However, there exists a significant gap between multi-modal and text-only contexts. In order to construct effective and reliable reasoning paths in multi-modal contexts, we characterize VoCoT with two features: (1) Object-centric. Objects are the basic semantic units in images and can serve as anchors to establish connections between multi-modal contextual information. Therefore, VoCoTs are required to include important objects, followed by relevant information extraction and analysis. (2) Visually-grounded. Key objects included in VoCoT should be represented by tuples of <text description, coordinates, visual object representation>. During inference, LMMs are required to generate both text and coordinates for objects to explicitly ground them within the images. The visual representation of objects further enhances the cross-modal relevance in the reasoning paths. Section 3.1 introduces how to obtain the visual object representations within current LMM frameworks.
2.2 VoCoT-Instruct-80K Dataset
The community has witnessed a surge in multi-modal instruction-following datasets Liu et al. (2024b); Luo et al. (2023); Chen et al. (2024). However, none of these datasets meet the requirements of the VoCoT format, which includes responses to instructions (1) with CoT-formatted multi-step reasoning processes and (2) with visually grounded object-centric information, i.e., objects with corresponding coordinates. In this section, we introduce the pipeline to construct a VoCoT-formatted dataset from three types of existing data sources.
Type 1: GQA Source
GQA Hudson and Manning (2019) is a VQA dataset that includes structured information: each image is paired with a scene graph, and a SQL-like reasoning path over the scene graph is provided for each VQA pair. An example is shown by the first part of Table 6 in Appendix A.1. Inspired by Shikra Chen et al. (2023b), we use a rule-based method to verbalize the SQL-like statements “[SementicStr]" and answers “[FullAnswer]" into fluent textual thoughts, supplementing objects descriptions with the corresponding coordinates from the scene graph.
Type 2: VQA-Based Source
Another intuitive way to construct data in VoCoT format is to supplement VQA data with multi-step reasoning processes in the middle of the Q2A process. With the assistance of GPT-4V, reasoning thoughts are generated based on images, questions, answers, and object information within the images. The second part of Table 6 provides an example. Furthermore, we control the output format through in-context learning. Specifically, a crafted sample is included in the input context. As overly simple questions may not require complex reasoning, we sample a subset of data from complex reasoning problems in LLaVA-Instruct (Liu et al., 2024a) as the source.
Type 3: Image-Only Source
Although the aforementioned two construction methods are effective, the generated data is limited to existing questions. To enhance the richness of questions and reasoning logic, we leverage GPT-4V to expand the constructed dataset. As illustrated by Table 6 in Appendix A.1, GPT-4V is provided with images and object information and prompted to generate complex questions, along with VoCoT-formatted reasoning paths and answers. In-context samples are also incorporated to ensure the correct output format. We choose LVIS Gupta et al. (2019) as the data source due to the diversity of objects included.
Ultimately, we construct VoCoT-Instruct-80K, comprising 72K, 6K, and 2K samples from data sources of Type 1, 2, and 3, respectively. More details about the construction process, including the rule-based conversion approach, prompts for GPT-4V, in-context samples, and quality control methods used are provided in Appendix A.1.
Figure 2: Illustration of the VolCano framework. The input and output are shown below and above the model, respectively. The blue and green rounded rectangles represent textual and visual tokens, respectively. Special tokens “[c]” and “[/c]” denotes the beginning and end of the coordinates (“[coor.]” in the figure). Coordinates are represented in text. In the output, we visualize coordinates by drawing corresponding boxes in the image for a better illustration. RefBind obtains the representations of objects with the image features and predicted coordinates.
Figure 3: Illustration of the RefBind mechanism.
3 VolCano: A VoCoT-enhanced LMM
In this section, we introduce how to adapt a modern LMM to utilize the VoCoT framework. We present the architecture of VolCano in Section 3.1, and detail the model training process in Section 3.2.
3.1 Architecture
As presented in Figure 2, the overall architecture of VolCano mainly follows LLaVA Liu et al. (2023b). VolCano is built on top of a decoder-only LLM as the backbone. We incorporate a vision transformer (ViT) as the visual encoder to encode image inputs. A two-layer MLP is adopted as the connection module to map the output of the visual encoder into the input space of the language backbone.
Representations of Multi-modal Sequences
VolCano represents image-text data as an interleaved sequence of visual and textual tokens. Text inputs are tokenized and represented using the embedding layer. Images and objects can appear at any position in the sequence and are represented by visual tokens. Images are encoded by the ViT. The obtained 2D feature maps are flattened into 1D sequences and further mapped to visual input tokens through the connection module.
Following the configuration of VoCoT, each object is represented by a visually grounded format: “{textual description} [c] {coordinates} [/c] {visual representation}”, e.g., “dog [c] 0.27, 0.08, 0.92, 0.81 [/c] $V_{\textrm{dog}}$”. “[c]” and “[/c]” are special tokens denoting the beginning and ending of coordinates. We use bounding boxes $[x_{\textrm{min}},y_{\textrm{min}},x_{\textrm{max}},y_{\textrm{max}}]$ as coordinates, $x$ and $y$ are normalized between 0 and 1 w.r.t to the image size. Coordinates are treated as text, undergoing tokenization and embedding.
In addition to text and coordinates, visual tokens of objects, such as $V_{\textrm{dog}}$, are supplemented to help the model reference the corresponding visual information in images. The visual tokens of objects are obtained based on the coordinates and image tokens through the RefBind mechanism. Once the end of coordinates token, “[/c]”, is detected in the input or generated, RefBind is activated to obtain the visual object tokens based on the coordinate between “[c]” and “[/c]”. The obtained object tokens are further appended after the “[/c]” token.
RefBind
A straightforward method for representing objects is to crop the corresponding regions in the image and encode them with the ViT. However, this method introduces additional computational costs and loses the contextual information of the complete image. For regions with very few pixels, representing objects with sub-images would introduce redundant information. To tackle with above issues, we propose the RefBind mechanism.
RefBind (short for “Reffering Bind”) is conceptually illustrated in Figure 3. Inspired by the RoI-pooling method in Fast-RCNN Girshick (2015), given a bounding box and the encoded 2D grid features of the entire image, we can efficiently index the patches in which the target object appears. The features of these patches are flattened into a sequence that represents the object. RefBind relies solely on indexing operations without additional computation. Additionally, the object representation obtained by RefBind inherently preserves contextual information within the whole image.
3.2 Training
The training of VolCano undergoes three stages:
Stage 1: Alignment Pre-training
The first stage aims to align visual representations with the LLM backbone. We utilize image-caption pairs from LLaVA-Pretrain Liu et al. (2023b). Only the parameters in the connection module are updated.
Stage 2: Multi-modal Interleaved Pre-training
Following the alignment stage, the model is trained to adapt to multi-modal interleaved sequences and visually grounded object representations with RefBind. Three types of sequences are considered: (1) Image-caption pair constitutes the simplest form of multi-modal sequences. We utilize ALLaVA-Caption Chen et al. (2024) which provides detailed descriptions. (2) Multi-modal document includes multiple image-text pairs in a sequence. Based on the relevance between images and text, we filter a subset of documents from MMC4 Zhu et al. (2024). (3) Grounded image caption further annotates the coordinates of objects in the caption. We extend the object representations to the visually grounded format consistent with VoCoT. Flickr30K Entities Plummer et al. (2015) and a subset of GRIT Peng et al. (2023) are adopted. Both the connection module and the LLM backbone are trained to model the multi-modal sequences.
Stage 3: Instruction Tuning
The pre-trained model is further fine-tuned to follow instructions in multi-modal contexts and perform multi-step reasoning with VoCoT. We supplement the constructed VoCoT-Instruct-80K and referring expression data Kazemzadeh et al. (2014); Chen et al. (2023b) to the existing non-CoT-form visual instruction data Liu et al. (2023b). We update the LLM backbone and connection module in this stage.
Table 1 summarizes the data mixtures used during the three training stages of VolCano. For more details, please refer to Section 4.1 and Appendix A.3.
4 Experiments
Stages Data Type Source Size
Stage 1 Image-Caption LLaVA 558k
Stage 2 Image-Caption ALLaVA 695k
Grounded Image-Caption GRIT 756k
Flickr30k 148k
Multimodal Document MMC4 890k
Stage 3 Visual Instruction LLaVA 612k
Referring Expression Shikra-RD 6k
RefCOCO 42k×$\times$3
g-RefCOCO 79k
VoCoT This Work 80k
Table 1: The training data mixtures used by VolCano.
4.1 Experiment Settings
Implementation Details
We build VolCano with the pre-trained ViT-L/14 CLIP Radford et al. (2021) visual encoder and Mistral-7B Jiang et al. (2023) as the baseline backbone. In addition, we explore the impact of a more powerful LLM backbone in our framework, constructing VolCanoQ2 based on Qwen2-7B Yang et al. (2024). Detailed parameter settings are provided in Appendix A.3. To save resources, we merely evaluate VolCanoQ2 in the main experiments and primarily focus on the Mistral-based VolCano in further analysis.
Evaluation Benchmarks
To validate the effectiveness and versatility of the VoCoT framework, we adopt different tasks across various scenarios for assessment: (1) General VQA benchmarks, including GQA Hudson and Manning (2019), MMBench Liu et al. (2023c), and SEED Li et al. (2023a); (2) Composite tasks requiring multi-step analysis and composite capabilities, such as visual spatial reasoning in VSR Liu et al. (2023a) and EmbSpatial Du et al. (2024), visual search in V-Star Wu and Xie (2023), complex reasoning in CLEVR Johnson et al. (2017) and Winoground Thrush et al. (2022), and complex referring expression in CLEVR-Ref Liu et al. (2019); (3) Hallucination benchmarks, including POPE Li et al. (2023d) and AMBER Wang et al. (2023b), to evaluate whether VoCoT can mitigate hallucinations. AMBER uses CHAIR Wang et al. (2023b) as the evaluation metric while accuracy is reported for other datasets. Details on the evaluation processes are provided in Appendix A.4.
Baselines
We compare VolCano with existing LMMs with ~7B parameters, as listed in Appendix A.5. For strict comparison, we construct a baseline model, VolCano-SE, which is based on the same architecture as VolCano but without VoCoT-Instruct-80K training data, so it can only perform single-step reasoning. We divide models into two groups for comparison: models based on baseline backbones (LLaMA-1,2, Vicuna, Qwen, and Mistral) and models based on advanced backbones (LLaMA-3 and Qwen2). We focus on models with single-image inputs in the main part. Please refer to Appendix B.1 for comparison and discussion involving models that use multiple additional sub-images as inputs for resolution enhancement.
Model General VQA Composite Tasks Hallucination
[…]
Models based on baseline LLM backbones
InstructBLIP-7B $224^{2}$ 1.3B 49.20 36.00 - 52.10 33.41 - 34.02 - - 72.10 8.80
Shikra-7B $224^{2}$ 0.3B - 58.80 - - 34.75 - - - - 83.10 -
mPLUG-Owl2-7B $448^{2}$ 0.3B 56.10 64.50 59.99 - 36.72 43.22 36.12 63.38 - - 10.60
MiniGPT-v2-7B $448^{2}$ 1.3B 60.10 55.14 51.50 62.90 43.85 46.23 33.19 62.00 24.90 80.50 -
Qwen-VL-Chat $448^{2}$ 1.9B 57.50 60.60 64.70 - 38.68 53.20 45.80 - 22.35 84.70 5.50
LLaVA1.5-7B $336^{2}$ 0.3B 62.00 64.30 53.80 64.24 42.43 43.73 48.68 55.31 6.70 84.50 7.80
VILA-7B $336^{2}$ 0.3B 62.30 61.50 60.40 66.02 38.05 47.60 46.22 66.37 - 84.50 10.50
VisCoT-7B $336^{2}$ 0.3B 63.00 63.82 63.23 - 37.01 53.15 61.76 56.40 32.05 86.10 7.20
VolCano-SE $336^{2}$ 0.3B 59.91 61.15 54.15 63.42 36.14 51.70 44.96 64.00 21.70 84.50 6.70
VolCano $336^{2}$ 0.3B 64.40 68.10 64.50 67.18 58.29 56.17 58.40 68.37 33.95 86.50 4.60
Models based on advanced LLM backbones
VILA1.5-8B $384^{2}$ 0.4B 63.50 64.38 64.41 53.76 54.95 55.22 58.74 66.00 - 84.90 8.50
Bunny-8B V1.0 $384^{2}$ 0.4B 64.00 70.86 67.59 65.71 53.54 54.47 58.32 68.50 - 86.40 7.40
VolCano-SEQ2 $336^{2}$ 0.3B 62.23 66.87 64.51 69.37 55.19 51.58 56.30 66.63 23.90 85.20 8.00
VolCanoQ2 $336^{2}$ 0.3B 64.60 71.61 66.95 74.22 59.86 56.78 62.81 68.78 34.00 86.60 4.40
GPT-4V $2048^{2}$ 75.80 71.60 68.24 36.07 51.90 55.00 83.75 - 82.00 4.60
[…]
Zero-Shot CoT < T > 56.79 52.47 51.70 45.32 57.20 65.00 67.50 52.20 6.70
Text CoT < T > 63.36 59.20 49.60 47.90 68.49 65.75 84.63 49.30 5.50
Coor. CoT < T, C > 64.32 58.59 54.42 53.78 66.86 65.87 85.47 47.80 4.30
Sub-Img CoT < T, C, S > 61.29 54.10 51.85 63.45 66.23 58.87 85.77 47.80 4.60
VoCoT < T, C, R > 64.50 58.29 56.17 58.40 67.18 68.37 86.50 51.00 4.60
Table 3: Comparison between different CoT formats. The “Obj-Format” column indicates the representation format of objects. T, C, S, and R are short for texts, coordinates, sub-images, and RefBind representations.
4.2 Main Results
Table 2 presents a thorough evaluation of existing LMMs. Several insights can be gleaned: (1) By comparing VolCano and VolCano-SE, it demonstrates that VoCoT effectively mitigates hallucinations and brings consistent improvement across all benchmarks. Section 4.3 delves deeper into how VoCoT contributes to reliable and visually grounded reasoning. (2) Across different datasets, VolCano and VolCanoQ2 achieve the best or second-best results within their respective group, where the advantages are more pronounced in composite tasks. On benchmarks like CLEVR and EmbSpatial, VolCano with a limited scale even outperforms powerful GPT-4V. (3) Furthermore, we compare two multi-modal CoT methods: VisCoT and VoCoT. VisCoT Shao et al. (2024) designs a simple two-step reasoning process: first searching for a single relevant region and then answering based on the detected region. Experimental results imply that VisCoT merely performs better on V-Star because the questions in V-Star perfectly align with the two-step search ligic of VisCoT. However, VisCoT falls short in other complex scenarios that involve interaction between multiple objects, indicating that VoCoT is a more generalizable format of multi-modal CoT. Overall, the experimental results validate the effectiveness of VoCoT-based multi-step reasoning in various scenarios. In addition, we find that VoCoT could seamlessly generalize to other scenarios including scene-text-centric tasks, please refer to the results and analysis in Appendix B.2.
Figure 4: Qualitative analysis to compare VoCoT and text-only CoT. Hallucinations are underlined.
[…]
Type 1 Type 2 Type 3
63.63 47.78 56.07 57.14 68.90 66.00 84.80 33.70 1.80
65.45 58.21 54.93 61.34 66.85 65.00 80.16 48.90 5.00
64.20 57.24 54.17 53.36 68.98 66.87 86.50 48.90 4.40
64.50 58.29 56.17 58.40 67.18 68.37 86.50 51.00 4.60
61.62 57.22 49.18 53.36 67.13 63.62 85.60 46.90 5.50
Table 4: Ablation of VoCoT-formatted data on stage 2 and stage 3 training process.
Analyzer VolCanoV VolCano VolCanoQ2 GPT-4V
Judger VolCanoV Vicuna-1.5 Mistral VolCano Vicuna-1.5 Mistral GPT-4
Accuracy(%) 63.5 64.5 67.2 67.2 65.1 67.8 73.8 74.2 68.2
Table 5: Performance with different analyzers and judgers on the VSR benchmark.
4.3 Comparing CoTs in Different Formats
We validate the effectiveness of the VoCoT format by comparing it with different CoT formats: (1) Zero-Shot CoT directly prompts VolCano-SE to think step-by-step without training; (2) Text CoT represents objects with only text descriptions; (3) Coor. CoT augments Text CoT with coordinates; and (4) Sub-Img CoT encodes sub-images as representations of objects rather than using RefBind.
Table 3 lists the results. Firstly, the zero-shot multi-step reasoning capability of VolCano-SE is limited. It is likely to exhibit hallucinations, highlighting the necessity to construct visual CoT tuning data. Secondly, CoT expressed only in texts is also affected by hallucinations, handling spatial reasoning well where each type of object appears only once, but failing to manage more complex scenarios. Thirdly, introducing coordinates grounds the thoughts to visual signals, mitigating hallucination and improving performance across various tasks. Furthermore, representations obtained by RefBind effectively help the model to utilize visual signals of objects. In contrast, the performance of Sub-Img CoT is overall inferior to that of Coor. CoT, which supports our claim in Section 3.1: simply encoding each object as an sub-image may introduce redundant information and degrade the performance.
Besides quantitative results, we present cases in Figure 4. We observe that text-only CoT is limited in terms of: (i) It may fail to accurately find/locate target object (Case 1). (ii) It is unable to leverage object-level visual information for inferring object attributes, as VoCoT does through RefBind (Case 3). (iii) It cannot resolve ambiguity between multiple objects (Case 4). (iv) Lack of interpretability. In general, it is crucial to ground the reasoning process to the visual information and VoCoT is the most suitable format.
4.4 Ablation on the Constructed Dataset
In Table 4, we explore the role of three types of data in VoCoT-Instruct-80K. The results implies that: (1) Type 1, the GQA-based data, is precise but limited in terms of diversity. Models trained solely on Type 1 data produces the fewest hallucinations but struggle to handle diverse questions. (2) Type 2 and 3 data effectively help the model generalize across various instructions. Nevertheless, totally removing Type 1 data will increase the risk of hallucinations. (3) Introducing multi-modal interleaved data in Stage 2 leads to a significant improvement. In summary, interleaved pre-training data and three types of VoCoT data should be jointly utilized.
4.5 Further Analysis
VoCoT enhances performance in complex questions
Figure 5 compares the performance of VolCano-SE and VolCano on questions of varying difficulty in CLEVR. The fitted curves and confidence intervals imply that as the number of required reasoning steps increases, the advantage of multi-step reasoning becomes more pronounced.
Figure 5: Relationship between performance and the number of reasoning steps required by the questions.
Disentangling Multi-Modal Reasoning
Our preliminary study finds that sometimes VolCano generates reasonable reasoning paths but fail to infer the correct answer. Therefore, we split the reasoning process into two sub-processes: analysis and judgment, where the former constructs reasoning paths and the latter provides conclusions. We conduct experiments to combine different analyzers and judgers on the VSR benchmark, where each object category corresponds to a single object in the image, allowing us to use text-only LLMs to judge based on the paths analyzed by VolCano. In Table 5, VolCanoV represents the Vicuna-based VolCano to explore the impact of LLM backbones. The results indicate that the judger plays a important role. The path analyzed by VolCano help GPT-4 make better decisions (73.8%) than GPT-4V (68.2%). However, the overall capability of VolCano is upper-bounded by the judgement ability of its LLM backbone. Comparison between VolCanoV, VolCano, and VolCanoQ2 further reveals the potential of applying VoCoT on stronger LLM backbones.
Case Study
Examples in Figure 6 show that VolCano provides a visually grounded description with no hallucinations in AMBER. In CLEVR, VolCano infers effective reasoning paths towards the answer. See Appendix B.8 for more cases in other datasets.
5 Related Works
5.1 Large Multi-Modal Models
Figure 6: Qualitative analysis with cases from AMBER and CLEVR. Hallucinations are highlighted.
Architecture of LMMs
A vast amount of research has emerged, focusing on adapting LLMs to handle multi-modal tasks. Initially, researchers treat LLMs as intelligent agents capable of using various tools. They train or prompt LLMs to invoke fundamental vision models, enabling them to complete multi-modal tasks such as captioning and VQA Wu et al. (2023); Yang et al. (2023c). Recent methods directly align the visual and textual representations in a unified backbone. To achieve this, various connection modules are designed, including MLP Liu et al. (2024b, 2023b), Q-Former Li et al. (2023b); Dai et al. (2023), and cross-attention layers Alayrac et al. (2022); Wang et al. (2023c).
Visual Instruction Data Construction
To enable LMMs to follow instructions in multi-modal contexts, a line of research has focused on converting existing academic datasets into instruction data Dai et al. (2023); Li et al. (2023c). In addition, researchers also use powerful tools like GPT-4V to assist in bootstrapping and generating high-quality data Zhu et al. (2023); Liu et al. (2024b). Further efforts are devoted towards improving both the quality and scale of the generated data Chen et al. (2024); Zhao et al. (2023); Wang et al. (2023a).
Visual Grounding in LMMs
Beyond text descriptions, some researchers further empower LMMs to understand and extract fine-grained visual information through visually grounded representations like coordinates Peng et al. (2023); Chen et al. (2023b), masks Yuan et al. (2024) and visual prompts Cai et al. (2024). However, most of these LMMs still rely on single-step and text-only reasoning, failing to effectively leverage fine-grained information during the reasoning process.
5.2 Multi-Step Reasoning
CoT in LLMs
Chain of Thoughts (CoT) is a series of prompting techniques designed to facilitate LLMs in addressing complex problems by guiding them through intermediate steps. CoT is first proposed through in-context learning Wei et al. (2022), followed by the introduction of zero-shot CoT Kojima et al. (2022), Auto-CoT Zhang et al. (2022) and self-consistency Wang et al. (2023d). Subsequently, CoT are extended to more complex formats Yao et al. (2024); Besta et al. (2024).
Visually Enhanced Reasoning
To address complex multi-modal problems, Shikra Chen et al. (2023b) initially explores the potential of applying CoT to specific tasks with LMMs. SoM Yang et al. (2023a) and Scaffolding Lei et al. (2024) respectively incorporate segmentation maps and dot grids in images to assist LMMs in reasoning, but such information can only be utilized by proprietary models like GPT-4V. The most related work is VisCoT Shao et al. (2024), which designs a two-step CoT: first searching for a relevant region and then answering based on the additional region information. This method is effective but cannot model complex multi-step reasoning. Overall, CoT has not been comprehensively explored in LMMs and there lack appropriate reasoning formats that could be generalized to various scenarios and tasks.
6 Conclusion
In this paper, we introduce VoCoT, a visually-grounded and object-centric chain of thoughts format to assist LMMs in multi-step reasoning. We also curate a VoCoT-formatted dataset from existing resources to train LMMs to learn reasoning with VoCoT. Building on this, we develop VolCano, a model capable of multi-step reasoning using the VoCoT format. Comprehensive experimental results demonstrate the effectiveness of our approach.
Limitations
Our work, as an early exploration of CoT techniques in large multi-modal models, is limited in the following aspects. (1) Currently, VoCoT is designed for single-image context and not applicable to multi-image inputs like videos and image sequences. Additional special tokens or marks can be introduced to extend VoCoT to a two-step grounding for multiple images. Each object is first grounded to a specific image and then localized to a region within that image. We will explore such mechanisms in our future work. (2) The construction of VoCoT-formatted dataset is limited by the cost of calling proprietary models and can not effectively scale up. In future work, we will explore methods to reduce the cost of data construction, including using smaller or open-source models, collecting and converting more finely annotated data (such as DocVQA) in a manner similar to GQA, and simulating and generating data based on specific needs, similar to CLEVR. (3) The presented VolCano model is currently limited with respect to 7B-sized models due to the lack of computational resources. As implied by the experimental results in Section 4.5, we hope to demonstrate the potential of applying VoCoT to larger and stronger backbones as explored in textual CoT techniques.
Ethical Statement
The presented VoCoT-Instruct-80K dataset is sourced from open-source datasets including GQA Hudson and Manning (2019), LLaVA-Instruct Liu et al. (2024b), and LVIS Gupta et al. (2019). We carefully follow the license to use these datasets and ensure that they are applicable for research purposes. The original datasets have been widely adopted by relevant researchers and ensure no risk of privacy leakage or harmful information. Furthermore, during the data collection and construction, we perform balanced sampling based on the distribution of object categories to alleviate distribution bias. As mentioned in Appendix A.1, we also conduct human-in-the-loop quality control to ensure the final dataset has correct information without ethical issues. Please refer to Appendix B.9 for the detailed discussion. Currently, the presented models and dataset focus on English, we hope to expand to other languages in the future. Our work and artifacts are designed with the principle of universality and fairness, without any preference for specific demographic groups.
Acknowledgment
The work is supported by National Key R&D Program of China (Grant Nos. 2023YFF1204800) and National Natural Science Foundation of China (Grant Nos. 62176058). The project’s computational resources are supported by CFFF platform of Fudan University.
Appendix A Supplementary Details
Configuration Alignment Multi-modal Interleaved Instruction Tuning
Visual Encoder OpenAI-CLIP ViT-L/14 OpenAI-CLIP ViT-L/14 OpenAI-CLIP ViT-L/14
Backbone Init Mistral-Chat-v0.2-7B Stage1 Stage2
Optimizer AdamW AdamW AdamW
Optimizer Hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e^{-6}$ $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e^{-6}$ $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e^{-6}$
Global batch size 256 128 128
Peak learning rate of LLM 1e-3 1e-5 1e-5
Learning rate schedule Cosine Cosine Cosine
Training Epochs 1 1 1
Warm-up ratio 0.03 0 0
Weight decay 0.0 0.0 0.0
Gradient clipping 1.0 1.0 1.0
Input image resolution 336 * 336 336 * 336 336 * 336
Input sequence to LLM 2048 3072 3072
Numerical precision bfloat16 bfloat16 bfloat16
GPU Usage 8 NVIDIA A100 8 NVIDIA A100 8 NVIDIA A100
Training Time 12h 48h 30h
Table 7: The detailed training hyper-parameters of VolCano. Except for the backbones for initialization, VolCanoQ2 follows the same hyper-parameters.
A.1 Data Construction Details
Construction Methods
As for the Type 1 data construction, we utilize a rule-based conversion method, the mapping rules used are listed in Table 16. In terms of Type 2 and Type 3 data, the prompts for GPT-4V are respectively shown in Table 17 and  18. We use in-context learning to let GPT-4V generate thought with multi-step reasoning path in VoCoT format.
Quality Control
For Type 1 data generated by rule based mapping, the quality is controled by the original source, namely GQA. We do not perform additional quality control.
For Type 2 and Type 3 data generated by GPT-4V, we first perform balanced sampling for images to achieve a balanced distribution of objects included. Secondly, we manually sample and check 200 data samples and find that initially constructed data suffers issues like uneven question types and incorrect object information. The first issue can be addressed by including well-designed in-context samples. The second issue is mainly caused by the potential incomplete and incorrect labels in LVIS. We find if GPT-4V think the information we provide is not enough to generate correct reasoning path, the response will contains error messages. All error message have patterns in common which may contains the following phrases: “From the object information provided", “provided object information", “From the bounding boxes provided", and so on. We remove the samples containing these patterns, leaving 8k samples after the filtering. Ultimately, by manually checking the constructed dataset, we do not observe any bias issues and achieve a 98% pass rate on the presented dataset.
A.2 Training Data Details
Stages Data Type Source Size
Stage 1 Image-Caption LLaVA-Pretrain Liu et al. (2024b, 2023b) 558k
Stage 2 Image-Caption ALLaVA-Caption Chen et al. (2024) 695k
Grounded Image Caption GRIT Peng et al. (2023) 756k
Flickr30k-Entities Plummer et al. (2015) 148k
Multimodal Document MMC4 Zhu et al. (2024) 890k
Stage 3 Visual Instruction LLaVA Liu et al. (2023b) 612k
Referring Expression Shikra-RD Chen et al. (2023b) 6k
RefCOCO Kazemzadeh et al. (2014) 42k
RefCOCO+ Kazemzadeh et al. (2014) 42k
RefCOCOg Kazemzadeh et al. (2014) 42k
g-RefCOCO He et al. (2023) 79k
VoCoT This Work 80k
Table 8: The data mixture used in the three training stages of VolCano.
We present our data mixture details in Table 8. In Stage 1, we use LLaVA-pretrain Liu et al. (2023b) dataset for projector alignment which contains 558k image caption pairs. In Stage 2, to better adapt to multi-modal interleaved sequences and visually grounded object representations with Refbind, we mix three types of data: (1) multimodal documents, (2) grounded image captions, and (3) high-quality image captions. Multimodal documents data is sample from MMC4 Zhu et al. (2024) by choosing which average similarity score between image and sentence before each image is larger than 0.3. Each multimodal document have multiple image, we remove samples with more than 6 images. Grounded image captions is from GRIT Peng et al. (2023) and Flickr30K Entities Plummer et al. (2015). For GRIT, we filter samples with clip score is larger than 0.35. We also use high quality image caption from ALLaVA Chen et al. (2024) which is generated by GPT-4V and provide details description. In Stage 3, we remove samples from LLaVA-Instruct that meet two criteria: sourced from RefCOCO, and sourced from VG where the object sub-image size is less 50, the reason is that we find extremely small regions in VG are probably with low quality.
A.3 Model & Training Details
The hyper-parameters in each stage are in Table 7. The learning rate setup mainly follows that of LLaVA Liu et al. (2024b). In stage 1, a large learning rate is used to update the connection module, aiming to quickly align the cross-modal representations. In the latter stages, a small learning rare is adopted to carefully fine-tune the backbone. Following Kosmos2 Peng et al. (2023), we introduce a special token “<grounding>” at the beginning of the sequence to control whether to require VolCano to produce visually grounded description.
Also notice that VolCano is able to perform both single-step and multi-step reasoning. Following Kojima et al. (2022), we introduce a prompt “Answer the question and include the reasoning proess. Locate key objects and provide bounding boxes in your thoughts.” as a trigger to tell the model whether to use VoCoT or not during the generation process.
A.4 Evaluation Details
In this section, we introduce the details in the evaluation procedure.
Category General VQA Spatial Reasoning Hallucination
Dataset GQA MMBench SEED VSR EmbSpat. POPE AMBER
Split testdev_balanced DEV Image test unseen test adversarial generative
Size 12578 4329 14233 1222 3625 3000 1004
Table 9: Information of Evaluation Benchmarks.
Category Composite Tasks Referring Expression
Dataset V-Star Wino CLEVR CLEVR-Ref
Split - test val -
Size 238 800 6000 2000
Table 10: Supplementation of Table 9.
A.4.1 Benchmark Details
General VQA Benchmarks
In GQA, we utilize the “testdev_balanced” split following Liu et al. (2024b). As for MMBench, we adopt the “DEV” split for the evaluation efficiency. In terms of SEED, we only consider the subset that the visual inputs are images. For GQA, we append a prompt “Please answer in a word or short phrase” to require models produce concise outputs.
Spatial Reasoning Benchmarks
For VSR, we utilize the unseen test split for zero-shot evaluation. For each sample, a description is provided and the model is required to distinguish if the claims is supported by the image. We use the prompt “Is there a event {description} in the image?” for this dataset. With respect to EmbSpatial, we use the test split for assessment.
Hallucination Benchmarks
For POPE, we consider the adversarial subset since it is the most challenging split. In AMBER, we leverage the generative task which asks the model to describe the image. All prompts are adopted from the original datasets with a yes-or-no instruction for POPE.
Benchmarks for Composite Tasks
For CLEVR, we utilize the val split. Because the original CLEVR validation set is too large, we categorize the data the into six types based on the question type: count, yes/no, shape, material, size, and color. We sample 1k questions from each category as test samples and construct a multiple-choice candidate set based on the feasible answers in the dataset. We will also open-source this subset. For Winoground, we utilize the test set and consider it as a caption selection multiple-choice question, the prompt is designed as “Please describe the image.”. As for V-Star, we directly use the V-Star benchmark. Regarding CLEVR-Ref, which is a referring expression task with relatively complex queries, we use the provided set for evaluation. We design a prompt as "Can you locate {phrase} in the image?" where “{phrase}” is the target query.
Please see Table 10 and Table 9 for the splits and scales of benchmarks used in this paper.
A.4.2 Evaluation Methods
All evaluation benchmarks adopted in this paper can be divided into three categories based on the task formulation: multiple-choice questions, open-ended generation, referring expression.
Multiple-Choice Question
For multiple-choice questions, we utilize the likelihood-based evaluation method, which is also known as the perplexity-based method. These methods are widely adopted in evaluating LMMs Li et al. (2023a, b); Dai et al. (2023); Li et al. (2023e). The key idea is to select the option with the highest generated likelihood, please refer to these papers for the detail. If VoCoT is utilized, the likelihood is computed based on the question, image, and the genrated reasoning path.
Open-Ended Generation
For GQA, we use the evaluation script provided by LLaVA Liu et al. (2023b) for a fair comparison. As for VSR and POPE, we require the model to answer in yes and no, enabling us to evaluate the correctness with exact match. With respoect to AMBER, we use the official evaluation method to assess the hallucinations in the generated descriptions.
Referring Expression
We first extract the predicted boxes from the outputs based on rules, then calculate the IoU between the ground truth box and the predicted box. If the IoU is larger than 0.5, it is considered as a correct prediction following  Kazemzadeh et al. (2014).
Further Analysis Setup
In the reasoning capability assessment part in Section 4.5, to leverage LLM as the judger model. We utilize a prompt “There is a image, {reasoning path}, please determine whether {description}, please answer yes or no.”, where the reasoning path are generated by the analyzer (with the coordinates and visual information removed), descriptions are the target description in VSR. If the model chooses not to predict, we consider the prediction as “no”.
A.5 Introduction to Baseline Models
We compare VolCano to several existing SOTA open-source LMMs, including BLIP-2 Li et al. (2023b), InstructBLIP Dai et al. (2023), Shikra Chen et al. (2023b), mPLUG-Owl2 Ye et al. (2023), MiniGPT-v2 Chen et al. (2023a), Qwen-VL-Chat Bai et al. (2023), VILA Lin et al. (2024), LLaVA-1.5 Liu et al. (2023b), and the most related VisCOT Shao et al. (2024). These models are based on baseline LLM backbones released in 2023, including LLaMA Touvron et al. (2023a), LLaMA-2 Touvron et al. (2023b), Vicuna Chiang et al. (2023), Mistral Jiang et al. (2023), and Qwen Bai et al. (2023). For models based on recently proposed advanced backbones like LLaMA-3 Dubey et al. (2024) and Qwen2 Yang et al. (2024), we compare VolCanoQ2 with Bunny He et al. (2024) and VILA-1.5 Lin et al. (2024). The models listed before take a single image as input, which is consistent with VolCano and VolCanoQ2 for a fair comparison. Additionally, we include another series of SOTA LMMs that enhance input resolution by splitting a single image into multiple sub-images: LLaVA-1.6 Liu et al. (2024a), Deepseek-VL Lu et al. (2024), and Monkey Li et al. (2023f). As for GPT-4V, examples in Figure 1 and all results are obtained by calling openai API using the “GPT-4V” model between 2024/05/18 to 2024/05/30.
We only consider zero-shot performance in Table 2, except for GQA. If a model has been trained on a specific evaluation benchmark, we do not report the corresponding evaluation results. For example, in Table 2, the result of VisCOT on VSR is omitted because it uses the corresponding training data. For certain models, including Bunny and VILA, due to the lack of clear evaluation details, we re-evaluate their performance in the same setting to make a fair comparison with VolCano.
Appendix B Supplementary Results and Discussion
B.1 Comparison between VolCano and High-Resolution models
Model General VQA Spatial Reasoning Composite Tasks Hallucination
[…]
Models with single-image inputs
VolCano-SE $336^{2}$ 0.3B 59.91 61.15 54.15 63.42 36.14 51.70 44.96 64.00 84.50 6.70
VolCano $336^{2}$ 0.3B 64.40 68.10 64.50 67.18 58.29 56.17 58.40 68.37 86.50 4.60
Models with multiple-image inputs
LLaVA1.6-7B $672^{2}$ 0.3B 64.20 68.40 66.15 66.86 56.82 50.35 58.80 64.88 86.90 -
LLaVA1.6-7Bm $672^{2}$ 0.3B 64.80 69.00 67.72 63.77 56.55 51.85 60.08 65.75 86.70 -
Deepseek-VL-7B $1024^{2}$ 0.4B - 71.32 70.40 67.51 41.77 48.77 62.18 64.88 85.77 -
Monkey $896^{2}$ 1.9B 60.70 61.95 67.58 62.93 32.91 46.33 67.23 68.63 82.57 -
GPT-4V $2048^{2}$ 75.80 71.60 68.24 26.65 51.90 55.00 83.75 82.00 4.60
Table 11: Comparison between VolCano and resolution-enhanced models. The notations follows Table 2. LLaVA1.6-7Bm represents the Mistral-based LLaVA1.6-7B.
Model Scene-Text-Oriented Tasks Know. Reasoning Ins. Following Captioning
TextVQAN AI2D ChartQA DocVQA MMMU MathVista MMVet MMT COCO NoCaps
LLaVA1.5-7B 46.1 43.0 14.9 2.9 28.4 26.7 30.5 45.0 94.5 95.6
VolCano-SE 45.1 45.3 19.0 4.6 29.3 27.8 32.1 44.1 81.0 90.6
VolCano 48.9 45.6 19.5 10.6 33.3 27.9 32.9 45.4 100.4 103.5
Table 12: Peformance on additional benchmarks. TextVQAN represents the TextVQA benchmark without providing reference OCR like in LLaVA-1.5 Liu et al. (2023b). “Know.” and “Ins.” are respectively short for knowledge and instruction. The best performance for each dataset are bolded.
In Table 2 we compare models with single-image and relatively low-resolution inputs. Comparing VolCano with LMMs that enhance input resolution by introducing multiple-image inputs, we observe that these methods primarily improve the performance in general VQA and V-Star, as V-Star provides high-resolution input images Wu and Xie (2023). However, in tasks that require complex reasoning, the improvement brought by higher resolutions becomes less significant. VolCano either exceeds or perform comparably with these models in such tasks, indicating the superiority of introducing multi-step reasoning over enriching the input information in these scenarios.
Notice that the RefBind mechanism introduced in Section 3 can be directly extended to fit multiple split sub-images by mapping the predicted coordinates to patches from different sub-images. We leave exploring combining these two vertical research directions—enhancing input resolution and introducing multi-step reasoning–as future work.
Figure 7: Cases on additional benchmarks included in Appendix B.2.
B.2 Performance in Additional Benchmarks
The benchmarks presented in the main text primarily focus on object-centric scenarios, which align with the design of VoCoT. Recently, a line of research develops LMMs to handle scene-text-oriented scenarios including document understanding and chart information extraction. In this section, we explore whether VoCoT can adapt to other scenarios by conducting experiments on additional benchmarks: TextVQA Singh et al. (2019), AI2D Kembhavi et al. (2016), ChartQA Masry et al. (2022), and DocVQA Mathew et al. (2021) for scene-text-oriented tasks; MMMU Yue et al. (2023) and MathVista Lu et al. (2023) for reasoning based on knowledge and scene texts; MMVet Yu et al. (2023) and MMT Ying et al. (2024) for instruction-following; COCO caption Veit et al. (2016) and NoCaps Agrawal et al. (2019). For image captioning, we report the CIDEr Vedantam et al. (2015) metric following Li et al. (2023b) while accuracy is provided for other tasks. For efficiency, we only consider LLaVA-1.5, VolCano-SE, and VolCano for a fair and straightforward comparison to validate the effect of VoCoT.
Scene-Text-Oriented Benchmarks
As shown in Table 12, we can find: (1) Quantitatively, VoCoT can generalize to text and chart-centric scenarios and improve the performance, even though VoCoT-Instruct-80K does not include similar data. (2) Qualitatively, we present several samples in Figure 7. VoCoT can treat text blocks and chart areas as "objects". By locating and analyzing the corresponding regions, VoCoT easily generalizes to such tasks. (3) The cross-domain improvements are very exciting, and the potential of VoCoT in such tasks can be further unleashed by enhancing input resolution and constructing text-centric VoCoT data (with fine-grained annotations in TextVQA, AI2D…), which we leave as our future work.
Method Avg. # Tokens Generated Avg. # Visual Tokens Added Avg. Inference Time per Query
VolCano-SE 2.2 0 0.11s
VolCano 115.9 326 0.74s
Table 13: Statistics of single-step and multi-step inference in VSR.
Model MMBench SEED MMMU MMT MMVet
LLaVA1.5-7B 64.3 53.8 28.4 45.0 30.5
VolCano-SE 61.1 54.2 29.3 45.1 32.1
VolCano 63.3 57.5 31.2 45.2 32.5
Table 14: Performance of differnt models with single-step inference settings.
Knowledge Reasoning and Instruction Following
According to results in Table 12, we observe that: (1) Although MMMU and MathVista focus on reasoning involving commonsense, knowledge, and mathematical deduction, the results indicate that the generalization ability of VoCoT can still improve the performance. As shown in Figure 7, the framework of VoCoT to locate and analyze local objects and regions can generalize to such tasks. (2) Improvements brought by VoCoT in the absence of similar training data demonstrate its potential in such scenario, inspiring us to incorporate both visually grounded information and conceptual knowledge in VoCoT framework in our future work. (3) In general instruction-following datasets, MMT and MMVet, it is shown that VoCoT does not hurt the instruction-following ability and brings improvement, which is consistent with the findings in SEED and MMBench mentioned in our main paper. See Appenidx B.3 for further exploration.
Image Captioning
Results in Table 12 imply that VoCoT helps VolCano to perceive accurate information and produce visually grounded descriptions, improving the quality of generated captions.
Generally, the results validates that VoCoT can generalize across various tasks and demonstrates potential for further enhancement.
B.3 Do VoCoT Damage the Original Capabilities of LMMs?
Another concern about whether the proposed VoCoT framework brings negative effects to the original capabilities of a LMM. Noticing that the trained VolCano can perform both single-step and multi-step reasoning as mentioned in Appendix A.3 and our paper aims at unleashing VoCoT reasoning ability of LMMs without affecting the original abilities. To validate that, we evaluate VolCano with the same single-step inference strategy as other compared baselines on general benchmarks in Table 14. It can be seen that VolCano maintains the original ability for conventional single-step inference, performing close to the two baselines in such settings and even surpassing them in some tasks. This indicates that the VoCoT framework does not hurt the original and fundamental capabilities of LMMs.
B.4 Grounding Capabilities of VolCano
Besides the reasoning capability analyzed in Section 4.5, another key capability to ensure the effectiveness of VoCoT is the grounding. However, the commonly adopted RefCOCO Kazemzadeh et al. (2014) dataset is widely used in training LMMs and not applicable for zero-shot evaluation. Besides, another problem with RefCOCO is that the query is relatively simple. To address this, we consider CLEVR-Ref in the main part because complex queries are considered. As a step further, we are interested in the grounding capability of LMMs during the generation process, but it is difficult to evaluate under this setting. Therefore, we conduct a preliminary exploration.
Specifically, we evaluate the performance of models to produce grounded captions: requiring models to annotate objects with coordinates while describing the images. 100 images are sampled from the LVIS Gupta et al. (2019) validation set for evaluation. Pairwise evaluation is performed to compare the grounded contents generated by two models. Given the image, ground-truth object information, and responses from 2 models, the judge, GPT-4V, will score 2 responses from multiple perspectives (including both content accuracy and coordinates accuracy) and determine the winner.
Model LLaVA1.5 LLaVA1.5 VolCano
Visual Input Image Image + Object Info. Image
CLEVR Acc. 43.73 45.70 56.17
Table 15: Performance of models in CLEVR with different visual inputs. Acc. and Info. are short for Accuracy and Information, respectively.
As GPT-4V itself can not generate precise coordinates, we conduct a sanity check whether GPT-4V can evaluate the relevance between two bounding boxes described in texts. for ease of testing, we ask GPT-4V to evaluate responses in RefCOCOg that include single coordinates, and measure the Pearson correlation coefficient between "coordinate accuracy" judged by GPT-4V and the actual IoU. The coefficient is 0.932, indicating that GPT-4V can accurately judge the matching degree between coordinates represented in text. For ease of testing, we ask GPT-4V to evaluate responses in RefCOCOg that include single coordinates, and measure the Pearson correlation coefficient between "coordinate accuracy" judged by GPT-4V and the actual IoU. The coefficient is 0.932, indicating that GPT-4V can accurately judge the matching degree between coordinates represented in text.
Generally, we believe that the current setup can, to some extent, reflects the grounding abilities of models during generation. Ultimately, we compare VolCano with Qwen-VL-Chat and MiniGPTv2. The win rates of VolCano against Qwen-VL-Chat and MiniGPTv2 are 76.5 and 82.0, respectively, indicating that VolCano can perform better in simultaneously locating and describing visual contents.
B.5 Computational Efficiency
VoCoT leads to additional computational overheads compared with traditional single-step reasoning: (1) RefBind only introduces indexing operations without float-point calculations. Additional cost is caused by the visual tokens added to the sequences that will be processed. (2) Multi-step reasoning leads to additional computation by requiring more tokens to be generated. Since precise calculation of the computation cost is challenging, we provide empirical statistics in Table 13.
Notice that additional computation is inevitable in CoT methods. However, with Flash Attention Dao et al. (2022) and KV Cache methods used during generation, we found the increase in token quantities does not lead to excessive inference time. In the future, we will follow text CoT papers to explore efficient decoding methods which help improve the efficiency of VoCoT.
B.6 Can Other Open-Source LMMs Directly Utilize Object Information?
Besides the traditional single-step reasoning paradigm, we wonder whether open-source LMMs like LLaVA can utilize the provided ground truth object information to perform grounded reasoning and enhance the improvement. We conduct a experiment on CLEVR where gold object coordinates exist in the dataset. According to Table 15, it is observed that LLaVA benefits from the information but can not utilize it effectively. In contrast, VolCano can perform localization, analysis and reasoning on its own, showing clear superiority.
B.7 Potential Language Bias in Spatial Reasoning
As presented in 3, text-only CoT method performs the best in VSR, we attribute this phenomenon to two reasons: (i) scenarios in VSR are relatively simple, and (ii) the text-only models can better leverage the language bias in spatial relationships as a shortcut. Firstly, each type of object appears only once in an image, with a one-to-one correspondence between the text and the object. so coordinates are not required to resolve ambiguity.
Secondly, as noted in Kamath et al. (2023), spatial reasoning datasets exhibit some language biases (e.g., a television is more likely to be on a table rather than under it). We find that such biases are more likely to be exploited by text-only CoT models, while VoCoT-based VolCano relies more on analyzed visual information.
We conduct a experiment: replacing the images in VSR with completely black images and using the original queries to ask whether the corresponding spatial relationships exist in the image: (1) The VoCoT-based VolCano predicts "no" for 99% of the samples (in line with expectations). (2) The text-only CoT-based model predicted "yes" for 26% of the samples, achieving a 54.1% accuracy rate among these predictions (better than random choice for the binary questions). (3) This phenomenon demonstrates that the text-only CoT-based model is more prone to being influenced by language bias, which provides it with a shortcut and additional advantage in VSR.
B.8 Case Study
Different from the black-box single-step reasoning paradigm, VolCano produces interpretable responses with the reasoning paths in text. Please see Figure 8910111213 for cases from representative datasets.
B.9 Discussion on Potential Social Impacts and Bias
We discuss potential issues and our solutions from the following perspectives:
  1. 1.
    Visual Bias: object categories in object detection datasets are unevenly distributed (mainly a long-tail distribution). To address the issue, we perform a balanced sampling of images based the included object categories. For the constructed dataset, we performed manual sampling and inspection and applied some filtering methods mentioned in Appendix A.1 to improve data quality. By checking the final constructed datasets, we did not observe any significant bias issues.
  2. 2.
    Misinformation: LMMs may produce erroneous information, namely hallucinations. The design of visually grounded representation in VoCoT aims to mitigate object hallucinations, and the experimental results validate the effectiveness.
  3. 3.
    Privacy issue: The images we adopt come from open-source and widely used datasets. Our construction method does not introduce additional privacy risks. Furthermore, we believe our object-centric method can be utilized to detect potential privacy issues in images. In future work, methods like RLHF will be used to guide VolCano to avoid detecting and analyzing data with potential privacy issues.
  4. 4.
    Beyond the above concerns. We utilize existing resources and there is no issue regarding potential personal information leakage and offensive content. The utilized tool, GPT-4V, also possess the capability to avoid generating offensive content. We manually checked the constructed dataset to ensure there is no such issues. We will continue to follow current responsible AI methods to monitor and alleviate our model and dataset for any biases or issues.
B.10 Discussion on the Use of Utilized and Presented Artifacts
In this work, we utilize existing artifacts including the data resources (GQA Hudson and Manning (2019), COCO Lin et al. (2014), and LVIS Gupta et al. (2019)), pre-trained models (CLIP Radford et al. (2021), Mistral Jiang et al. (2023), and Qwen2 Yang et al. (2024)), and existing datasets as listed in Table 1. All utilized artifacts are open-sourced to the research community. We carefully follow the license to use artifacts and ensure they are applicable for the research purpose. All utilized artifacts mainly focus on the English domain while Qwen2 and Mistral both possess multi-lingual capabilities. Please refer to the original resource for other information about the artifacts.
As for the artifacts we presented in this paper, including VoCoT-Instruct-80K and pre-trained VolCano and VolCanoQ2, we will release the data, code, and model weights to the community for research purpose. Our introduced artifacts are primarily designed for the English domain and will be extended to more languages. Our artifacts are designed with the principle of universality and fairness, without any preference for specific demographic groups.
B.11 Usage of AI Assistants
In this work, we mainly utilize GPT-4V as the AI assistants for preliminary exploration as in Figure 1, data transforming as in Section 2.2, and as a intelligent agent to judge the performance of models (Appendix B.4). Besides that, we utilize ChatGPT to help polish some parts of this paper.
Opeation Mapping Rule
relate: sub, relation, obj Check the {subject} that is {arg2} {object}.
same: attribute [obj1,obj2] The question ask if the two objects has same {attribute}.
Check if they have same {attribute}.
common: [obj1,obj2] The question ask the common attribute of the two objects.
different: attribute, [obj1, obj2] The question ask if the two objects has different {attribute}
and: [obj1, obj2] The question ask about ’and’ relation.
select: obj1 Find {obj1}
exist: ? obj1 It doesn’t exist. if obj1 is not in annotation else It exist
verify: attribute,value, obj1 Verify if the {attribute} of {obj1} is {value}.
or: [obj1, obj2] The question ask about ’or’ relation.
choose: obj1, attribute, value1, value2, obj2 Think {obj1}’s {attribute} is {value1} or {value2} of {obj2}.
choose: obj1, attribute, value1, value2, Think {obj1}’s {attribute} is {value1} or {value2}.
Table 16: Mapping rule for transferring SQL-like query statement to string in GQA Source Type Data construction.
messages = [ {"role":"system", "content": f"""You are an excellent generator of image QA reasoning processes based on question-answer pairs and object information represented by object bounding boxes (x_left_top, y_left_top, x_right_down, y_right_down).Your task is to generate reasoning process based on the questions and answers you are given. The reasoning process should include the reasoning path, relevant object bounding boxes, and inference clues, including but not limited to the object’s number, location, and your own background knowledge. The object in your reasoning path must annotate with object bounding box. The bounding box must come from the object information given by the user, please do not detect it yourself! ! ! ! Don’t mention object information directly, just annotate it with bounding boxes.When you refer to the information in prompt, the text should show that you did not know the answer in advance, but that you reasoned it out yourself. And don’t directly say that something doesn’t appear in the information provided. Don’t mention anything in the prompt in your reply, and don’t mention bounding boxes in the generated reasoning process.You will follow instructions to the best of your ability. Your response should follow the following format: "Thought":"""""}, [Uncaptioned image] {"role": "user","content": [{"type": "text","text": "[IMAGE1]:[Object Info]: Coffee1: [0.04, 0.25, 0.20, 0.28], Bean: [0.22 , 0.12 , 0.29 , 0.32 ], Vegetable: [0.52, 0.11, 0.29, 0.31 ], Coffee2: [0.76, 0.25, 0.21, 0.27 ], Yam: [0.69, 0.48, 0.27, 0.34 ], Burrito: [0.34, 0.40, 0.35, 0.44 ], Orange: [0.05, 0.50, 0.27, 0.34 ], Chopsticks: [0.61, 0.69, 0.22, 0.31] [question]: What is the food with the most Vitamin C in this image?" [answer]: Orange."},{"type": "image_url", "imag_url": {"url": f"data:image/jpeg;base64,{image1}"}}]}, {"role": "assistant", "content":”’"thought": "From the picture, you can see two cups of coffee [0.04, 0.25, 0.20, 0.28],[0.76, 0.25, 0.21, 0.27 ], one portion of bean[0.22 , 0.12 , 0.29 , 0.32 ], one portion of vegetables[0.34, 0.40, 0.35, 0.44 ], one portion of yam[0.69, 0.48, 0.27, 0.34 ], one burrito[0.34, 0.40, 0.35, 0.44 ], and a plate of oranges[0.05, 0.50, 0.27, 0.34 ]. Among these foods, bean[0.22 , 0.12 , 0.29 , 0.32 ] contains protein, yams[0.05, 0.50, 0.27, 0.34 ] and burrito[0.34, 0.40, 0.35, 0.44 ] are rich in starch, vegetables[0.34, 0.40, 0.35, 0.44 ], and oranges[0.05, 0.50, 0.27, 0.34 ] are foods that may contain vitamin C, but oranges[0.05, 0.50, 0.27, 0.34 ] have a higher vitamin C content, so oranges[0.05, 0.50, 0.27, 0.34 ] are foods that contain more vitamin C."”’},]
Table 17: Prompt and one in-context sample for VQA-Based Source Type Data generation.
messages = [ {"role":"system", "content": f"""You are an excellent image describer and question-answer generator based on the image and object information which is represented by object bounding box (x_left_top, y_left_top, x_right_down , y_right_down). You have three tasks in total. Your first task is to ask a complex question that requires close inspection of the image and strong reasoning ability to answer. Your second task is to answer the question you raised solely based on the given image. Your third task is to generate the reasoning thought. The reasoning thought should contain the reasoning path, relative object bounding box and inference clue, include but are not limited to the object numbers, location and background knowledge from yourself. The bounding box must come from the user given object information, Do not detect by yourself !!!! Do not mention the object infomation directly, just annotate with bounding box. When you ask questions, try to find the most valuable information in the picture to ask about, and ask a question that is relevant to that information. When you ask questions, do not involve violence, advertisement, possible invasion of privacy, or questions that may cause discomfort. Do not mention anything from the prompt in your response and Do not mention bounding box in your generated question. You will follow the instructions to the best of your ability. Your response should follow the following format: "question":"","answer":"","Thought":"""""}, [Uncaptioned image] {"role": "user","content": [{"type": "text","text": "[IMAGE1]:[Object Info]: Coffee1: [0.04, 0.25, 0.20, 0.28], Bean: [0.22 , 0.12 , 0.29 , 0.32 ], Vegetable: [0.52, 0.11, 0.29, 0.31 ], Coffee2: [0.76, 0.25, 0.21, 0.27 ], Yam: [0.69, 0.48, 0.27, 0.34 ], Burrito: [0.34, 0.40, 0.35, 0.44 ], Orange: [0.05, 0.50, 0.27, 0.34 ], Chopsticks: [0.61, 0.69, 0.22, 0.31]"},{"type": "image_url", "imag_url": {"url": f"data:image/jpeg;base64,{image1}"}}]}, {"role": "assistant", "content":”’"question":"What is the food with the most Vitamin C in this image?", "answer": "Orange.", "thought": "From the picture, you can see two cups of coffee [0.04, 0.25, 0.20, 0.28],[0.76, 0.25, 0.21, 0.27 ], one portion of bean[0.22 , 0.12 , 0.29 , 0.32 ], one portion of vegetables[0.34, 0.40, 0.35, 0.44 ], one portion of yam[0.69, 0.48, 0.27, 0.34 ], one burrito[0.34, 0.40, 0.35, 0.44 ], and a plate of oranges[0.05, 0.50, 0.27, 0.34 ]. Among these foods, bean[0.22 , 0.12 , 0.29 , 0.32 ] contains protein, yams[0.05, 0.50, 0.27, 0.34 ] and burrito[0.34, 0.40, 0.35, 0.44 ] are rich in starch, vegetables[0.34, 0.40, 0.35, 0.44 ], and oranges[0.05, 0.50, 0.27, 0.34 ] are foods that may contain vitamin C, but oranges[0.05, 0.50, 0.27, 0.34 ] have a higher vitamin C content, so oranges[0.05, 0.50, 0.27, 0.34 ] are foods that contain more vitamin C."”’},]
Table 18: Prompt and one in-context sample for Image-Only Source Type Data generation.
Figure 8: Cases on CLEVR dataset.
Figure 9: Cases on GQA dataset.
Figure 10: Cases on Seed-Bench dataset. Seed-Bench provide a question and a list of answer options for each sample.
Figure 11: More Cases on Seed-Bench dataset.
Figure 12: Cases on VSR dataset. VSR dataset provides a statement for each image. The task is to judge whether the statement is right about this image.
Figure 13: Cases from the AMBER dataset. The task is to describe each image. We both present our VolCano and VolCano-SE responses for these cases. VolCano-SE is trained without VoCoT data. The underline phrase is hallucination generated by VolCano-SE.